### v1

# Partial Rewriting and Value Interpretation of Logically Constrained Terms

Abstract
Logically constrained term rewrite systems (LCTRSs) are a rewriting formalism that naturally supports built-in data structures, including integers and bit-vectors. The recent framework of existentially constrained terms and most general constrained rewriting on them (Takahata et al., 2025) has many advantages over the original approach of rewriting constrained terms. In this paper, we introduce partial constrained rewriting, a variant of rewriting existentially constrained terms whose underlying idea has already appeared implicitly in previous analyses and applications of LCTRSs. We examine the differences between these two notions of constrained rewriting. First, we establish a direct correspondence between them, leveraging subsumption and equivalence of constrained terms where appropriate. Then we give characterizations of each of them, using the interpretation of existentially constrained terms by instantiation. We further introduce the novel notion of value interpretation, that highlights subtle differences between partial and most general rewriting.
keywords:
existentially constrained term, most general constrained rewriting, partial constrained rewriting, interpretation of constrained terms, logically constrained term rewrite system
category:
1 Introduction
Logically constrained term rewrite systems (LCTRSs) are a formalism that extends term rewrite systems (TRSs). Many real-world computational problems involve built-in data structures, such as integers and bit-vectors, that often limit the applicability of techniques developed for standard TRSs. By equipping rewrite rules with constraints interpreted over an arbitrary model, the LCTRS formalism naturally supports such built-in data structures. A constrained rewrite rule can be applied only when its instantiated constraint is valid in the underlying model; consequently, the framework is sufficiently general to accommodate arbitrary built-in data structures in theory. Moreover, by incorporating support from SMT-solvers and leveraging the theories in common SMT libraries, LCTRS tools can effectively address rewriting problems involving these data structures.
Many topics from the term rewriting literature have already been studied within the LCTRSs formalism, including such as confluence analysis [14, 16], (non-)termination analysis [7, 11], completion [19], rewriting induction [9, 4], algebraic semantics [1], and complexity analysis [20]. Nevertheless, the framework also exhibits aspects that are not covered by traditional term rewriting. One of such aspect is the rewriting of constrained terms. A constrained term consists of a term together with a constraint that restricts the ways in which the term may be instantiated. For example, the constrained term $\mathsf{f}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]$ (in LCTRS notation) can be intuitively viewed as a set of terms $\{\hskip 0.85358pt\mathsf{f}(x)\mid x>2\hskip 0.85358pt\}$. In the analysis of LCTRSs, constrained rewrite rules must be handled, and consequently rewriting constrained terms naturally arises. Indeed, constrained rewriting is an integral part of many different analysis techniques. For instance, in confluence analysis, finding a specific joining sequence for constrained critical pairs often requires reasoning about two terms under a shared constraint, which leads to working on constrained terms. Similarly, in rewriting induction, rewriting of constrained terms is employed in several of its inference steps.
Rewriting constrained terms was already introduced in early work on LCTRSs [8]. However, the original formulation of rewriting constrained terms is difficult to deal with, since rewrite steps are used modulo equivalence transformations of constrained terms, and the results of rewriting a constrained term remain highly flexible even when both the rewrite rule and the position are fixed. These difficulties motivated us to revisit the formalism of constrained rewriting in LCTRSs in our recent paper [takahataPPDP]. In that work, we have introduced the notions of existentially constrained terms and most general constrained rewriting for left-linear LCTRSs. We have shown the uniqueness of reducts in this new setting, as well as commutativity between rewrite steps and equivalence, under the restriction to left-value-free rewrite rules [^1] However, any rewrite rule is simulated by a left-value-free rewrite rule, and thus this is not a real restriction.. Our formalism of most general constrained rewriting extracts the so-called “most general” part of the corresponding original formalism. As a result, it offers several advantages over the original approach of rewriting constrained terms.
In this paper, we introduce another notion of rewriting on existentially constrained terms, called partial constrained rewriting. The difference between the two notions can be explained informally as follows. On the one hand, most general constrained rewriting applies a rule to all instantiations of a constrained terms. On the other hand, partial constrained rewriting applies a rule to only some instantiations of a constrained term (see LABEL:fig:intuition_for_twokinds_of_rewriting).
Example 1.1.
[…]
$\begin{array}[]{lclc}\mathsf{sum}(x)\penalty 10000\ [1\leqslant x\land x\leqslant 5]&\to_{\mathcal{R}}&x+\mathsf{sum}(x+-1)\penalty 10000\ [1\leqslant x\land x\leqslant 5]\\ &\to_{\mathcal{R}}&x+\mathsf{sum}(y)\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1]$$\not\to_{\mathcal{R}}\end{array}$
obtains a normal form of most general constrained rewriting. On the other hand, we have $\mathsf{sum}(1)\to_{\mathcal{R}}^{*}1$, $\mathsf{sum}(2)\to_{\mathcal{R}}^{*}3$, $\mathsf{sum}(3)\to_{\mathcal{R}}^{*}6$, $\mathsf{sum}(4)\to_{\mathcal{R}}^{*}10$, $\mathsf{sum}(5)\to_{\mathcal{R}}^{*}15$. However, most general rewriting is not suitable for this computation, as there is no single rule that can be applied to the constrained term $\mathsf{sum}(y)\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1]$. In contrast, partial constrained rewriting allows the rewriting process to proceed by
[…]
because the first rule is applicable as the constraint $1\leqslant x\land x\leqslant 5\land y=x-1$ is valid with $x=1$ and $y=0$. Similarly, partial constrained rewriting allows us to continue using the second rewrite rule as
$\begin{array}[]{lcl}\cdots&\to_{\mathcal{R}}&1+\mathsf{sum}(y)\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1]\\ &\leadsto_{\mathcal{R}}&1+(1+\mathsf{sum}(z))\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1\land z=y-1]\end{array}$
Thus, by partial rewriting we obtain $\{t\in\mathsf{NF}\mid\mathsf{sum}(x)\penalty 10000\ [1\leqslant x\land x\leqslant 5]\leadsto^{*}t\}=\{1,3,6,10,15\}$.
The motivation for introducing this variant of rewriting on existentially constrained terms is twofold: not only is the idea itself natural, but it has also appeared implicitly in previous analyses and applications of LCTRSs, although the two rewriting formalisms may need to be applied separately depending on the case.
In this paper, we study the differences between most general rewriting and partial rewriting from several perspectives. First, we establish a direct correspondence between the two formalisms, showing that partial constrained rewrite steps are included in most general rewrite steps, but not vice versa. We furhter demonstrate that most general rewrite steps can be simulated by partial rewriting through subsumption and equivalence of constrained terms. We also provide characterizations of each type of rewriting constrained terms in terms of interpretations of constrained terms and rewrite rules as sets of terms and rewrite rules, respectively. The idea of interpreting (original) constrained terms as sets of terms has appeared, for example, in [4, Section 2.3] to reduce certain problems on LCTRSs to problems in TRSs. However, these characterizations alone are insufficient to distinguish partial constrained reductions from most general constrained reductions. This motivates the introduction of a novel notion of value interpretation. Like the standard interpretation, the value interpretation reflects the equivalence of constrained terms; however, unlike the standard interpretation, it does not reflect subsumption. Using the value interpretation, we provide more precise characterizations of the two notions of rewriting constrained terms, highlighting subtle differences between partial and most general rewriting.
Two kinds of rewrite steps by the ruleMost General Rewrite StepPartial Rewrite Step
Figure 1: Most General and Partial Rewrite Steps
The remainder of the paper is organized as follows. Section 2 introduces some preliminaries. In Section 3, we introduce partial constrained rewriting and establish direct correspondences between partial constrained rewriting and most general constrained rewriting. Section 4 explores properties of interpretations of existentially constrained terms, introduces the interpretation of constrained rewrite rules, and characterizes both partial and most general rewrite steps based on these interpretations, while also highlighting the limitations of these. In Section 5, we introduce the value interpretation and discuss its properties. Section 6 provides characterizations of partial and most general rewrite steps using value interpretations. Related work is discussed in Section 7, and Section 8 concludes the paper.
2 Preliminaries
In this section, we briefly recall the basic notions of LCTRSs [1, 8, 14, 16, 17] and fix additional notations used throughout this paper. Familiarity with the basic notions of term rewriting is assumed (see, e.g., [2]).
We employ a sorted signature $\langle\mathcal{S},\mathcal{F}\rangle$ consisting of a set $\mathcal{S}$ of sorts and a set $\mathcal{F}$ of function symbols, where each $f\in\mathcal{F}$ is attached with its sort declaration $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}$ ($\tau_{0},\ldots,\tau_{n}\in\mathcal{S}$). The set of sorted variables is denoted by $\mathcal{V}$. For $X\subseteq\mathcal{V}$ and $F\subseteq\mathcal{F}$, the set of term over the function symbols in $F$ and variables in $X$ is denoted by $\mathcal{T}(F,X)$. We write $t^{\tau}$ to denote that a term $t$ has sort $\tau$. The set of variables appearing in a term $t$ is denoted by $\mathcal{V}\mathsf{ar}(t)$. A term $t$ is said to be linear when any variable occurs at most once in $t$. For brevity, we write a sequence of variables as $\vec{x}$, etc. The set of variables occurring in $\vec{x}$ is denoted by $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$. Likewise, the set of sequences of elements of a set $T$ is denoted by $T^{*}$.
The set of positions in a term $t$ is denoted by $\mathcal{P}\mathsf{os}(t)$. The symbol and subterm occurring at a position $p\in\mathcal{P}\mathsf{os}(t)$ is denoted by $t(p)$ and $t|_{p}$, respectively. A term obtained from $t$ by replacing subterms at positions ${p_{1}},\dots,{p_{n}}$ by the terms ${t_{1}},\dots,{t_{n}}$, having the same sort as $t|_{p_{1}},\ldots,t|_{p_{n}}$, as $t[{t_{1}},\dots,{t_{n}}]_{{p_{1}},\dots,{p_{n}}}$ or just $t[{t_{1}},\dots,{t_{n}}]$ when no confusions arises. Sometimes we consider an expression obtained by replacing those subterms $t|_{p_{1}},\ldots,t|_{p_{n}}$ in $t$ by holes of the same sorts, which is called a multihole context and denoted by $t[\penalty 10000\ ]_{{p_{1}},\dots,{p_{n}}}$. A sort-preserving function $\sigma$ from $\mathcal{V}$ to $\mathcal{T}(\mathcal{F},\mathcal{V})$ whose domain $\mathcal{D}\mathsf{om}(\sigma)=\{\hskip 0.85358ptx\in\mathcal{V}\mid\sigma(x)\neq x\hskip 0.85358pt\}$ is finite is called a substitution, where it is identified with its homomorphic extension $\sigma\colon\mathcal{T}(\mathcal{F},\mathcal{V})\to\mathcal{T}(\mathcal{F},\mathcal{V})$. For a set of terms $U$, we write $\sigma(U)=\{\hskip 0.85358pt\sigma(t)\mid t\in U\hskip 0.85358pt\}$. As usual, $\sigma(t)$ is also written as $t\sigma$. We write $\{x_{1}:=t_{1},\ldots,x_{n}:=t_{n}\}$ for the substitution $\sigma$ such that $\sigma(x_{i})=t_{i}$ ($1\leqslant i\leqslant n)$ and $\mathcal{D}\mathsf{om}(\sigma)\subseteq\{x_{1},\ldots,x_{n}\}$, which also be written as $\{\vec{x}:=\vec{t}\}$ for brevity.
We assume that the set $\mathcal{S}$ of sorts and the set $\mathcal{F}$ of function symbols are partitioned into two disjoint sets, respectively: $\mathcal{S}=\mathcal{S}_{\mathsf{th}}\uplus\mathcal{S}_{\mathsf{te}}$ and $\mathcal{F}=\mathcal{F}_{\mathsf{th}}\uplus\mathcal{F}_{\mathsf{te}}$, where each $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}\in\mathcal{F}_{\mathsf{th}}$ satisfies $\tau_{i}\in\mathcal{S}_{\mathsf{th}}$ for all $0\leqslant i\leqslant n$. Elements of $\mathcal{S}_{\mathsf{th}}$ ($\mathcal{F}_{\mathsf{th}}$) and $\mathcal{S}_{\mathsf{te}}$ ($\mathcal{F}_{\mathsf{te}}$) are called theory sorts (symbols) and term sorts (symbols). Moreover, we consider a model $\mathcal{M}=\langle\mathcal{I},\mathcal{J}\rangle$ over the sorted signature $\langle\mathcal{S}_{\mathsf{th}},\mathcal{F}_{\mathsf{th}}\rangle$. Here, $\mathcal{I}$ is interpretations for theory sorts, and $\mathcal{I}$ assigns any sort $\tau\in\mathcal{S}_{\mathsf{th}}$ a non-empty set $\mathcal{I}(\tau)$; $\mathcal{J}$ is interpretations for theory symbols, and $\mathcal{J}$ assigns any theory symbol $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}\in\mathcal{F}_{\mathsf{th}}$ to a function $\mathcal{J}(f)\colon\mathcal{I}(\tau_{1})\times\cdots\times\mathcal{I}(\tau_{n})\to\mathcal{I}(\tau_{0})$. A valuation $\rho$ on the model $\mathcal{M}=\langle\mathcal{I},\mathcal{J}\rangle$ is a mapping that assigns any $x^{\tau}\in\mathcal{V}$ with $\rho(x)\in\mathcal{I}(\tau)$. The interpretation of a term $t$ in the model $\mathcal{M}$ over the valuation $\rho$ is denoted by ${t}]\!]_{\mathcal{M},\rho}$.
We further assume that any element $a\in\mathcal{I}(\tau)$ appears as a constant $a^{\tau}$ in $\mathcal{F}_{\mathsf{th}}$, for every $\tau\in\mathcal{S}_{\mathsf{th}}$. These constants are called values and the set of all values is denoted by $\mathcal{V}\mathsf{al}$. The set of values contained in a term $t$ is denoted by $\mathcal{V}\mathsf{al}(t)$. A term $t$ is value-free when $\mathcal{V}\mathsf{al}(t)=\varnothing$. For a substitution $\gamma$, we define $\mathcal{V}\mathcal{D}\mathsf{om}(\gamma)=\{\hskip 0.85358ptx\in\mathcal{V}\mid\gamma(x)\in\mathcal{V}\mathsf{al}\hskip 0.85358pt\}$.
We assume a special sort $\mathsf{Bool}\in\mathcal{S}_{\mathsf{th}}$ of boolean having the standard interpretation $\mathcal{I}(\mathsf{Bool})=\mathbb{B}=\{\hskip 0.85358pt\mathsf{true},\mathsf{false}\hskip 0.85358pt\}$, and the existence of necessary standard theory symbols such as $\neg$, $\land$}, $\Rightarrow$}, $\Leftrightarrow$}, $_{\tau}$}, etc. with the standard interpretations. The terms in $\mathcal{T}(\mathcal{F}_{\mathsf{th}},\mathcal{V})$ of sort $\mathsf{Bool}$ are called logical constraints (simply constraints). For a constraint $\varphi$ and a valuation $\rho$ over the model $\mathcal{M}$, we write $\vDash_{\mathcal{M},\rho}\varphi$ if ${\varphi}]\!]_{\mathcal{M},\rho}=\mathsf{true}$, and $\vDash_{\mathcal{M}}\varphi$ if $\vDash_{\mathcal{M},\rho}\varphi$ for all valuations $\rho$. The constraint $\varphi$ is said to be valid in this case. A constraint $\varphi$ is said to be satisfiable if $\vDash_{\mathcal{M},\rho}\varphi$ for some valuation $\rho$. For $X\subseteq\mathcal{V}$, a substitution $\gamma$ is said to be $X$-valued if $\gamma(X)\subseteq\mathcal{V}\mathsf{al}$. We write $\gamma\vDash_{\mathcal{M}}\varphi$ (and say $\gamma$ respects $\varphi$) if the substitution $\gamma$ is $\mathcal{V}\mathsf{ar}(\varphi)$-valued and $\vDash_{\mathcal{M}}\varphi\gamma$. If no confusion arises then we drop the subscript $\mathcal{M}$ in these notations.
For example, for the LCTRS as in Example˜1.1 we prepare $\mathcal{S}_{\mathsf{th}}=\{\mathsf{Bool},\mathsf{Int}\}$, $\mathcal{S}_{\mathsf{te}}=\varnothing$, $\mathcal{F}_{\mathsf{th}}=\mathbb{Z}\cup\mathbb{B}\cup\{+:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Int},{\geqslant}:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Bool},{<}:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Bool},\ldots\}$, and $\mathcal{F}_{\mathsf{te}}=\{\mathsf{sum}:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Int}\}$, where $\mathcal{V}\mathsf{al}=\mathbb{Z}\cup\mathbb{B}$. We set $\mathcal{I}(\mathsf{Bool})=\mathbb{B}$, $\mathcal{I}(\mathsf{Int})=\mathbb{Z}$, $\mathcal{J}(+):\mathbb{Z}\times\mathbb{Z}\to\mathbb{Z}$ is the addition function on $\mathbb{Z}$, etc. If we take a valuation $\rho$ such that $\rho(x)=\rho(y)=1$, then we have ${x+y}]\!]_{\mathcal{M},\rho}=2$ and ${x+y\geqslant 2}]\!]_{\mathcal{M},\rho}=\mathsf{true}$. Thus, $\vDash_{\mathcal{M},\rho}x+y\geqslant 2$, as $x+y\geqslant 2$ is a constraint. The constraint $x+y\geqslant 2$ is satisfiable but not valid. For example, $x=_{\mathbb{Z}}x$, $(x\geqslant 0)\lor(x<0)$ and $(0<x)\Rightarrow(x\geqslant 0)$ are valid constraints; we have e.g. $\vDash_{\mathcal{M}}(x\geqslant 0)\lor(x<0)$. Note that values are terms, and thus $\sigma=\{x:=1,y:=1,z:=\mathsf{sum}(w)\}$ is a substitution. We have $(x+z)\sigma=1+\mathsf{sum}(w)$. As $\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=\{x,y\}$, the substitution $\sigma$ is $\{x,y\}$-valued. We also have $\sigma\vDash_{\mathcal{M}}x+y\geqslant 2$.
An existential constraint is a pair $\langle\vec{x},\varphi\rangle$ of a sequence of variables $\vec{x}$ and a constraint $\varphi$, written as $\exists\vec{x}.\ \varphi$, such that $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\subseteq\mathcal{V}\mathsf{ar}(\varphi)$. We define the sets of free and bound variables of $\exists\vec{x}.\ \varphi$ as follows: $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)=\mathcal{V}\mathsf{ar}(\varphi)\setminus\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$ and $\mathcal{B}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)=\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$. We define the interpretation of existential constraint like this: $\vDash_{\mathcal{M},\rho}\exists\vec{x}.\ \varphi$ if there exists a valuation $\rho^{\prime}=\rho[\vec{x}\mapsto\vec{w}]$ which updates $\rho$ on $\vec{x}$ with $\vec{w}$ (i.e. $\rho^{\prime}(x)=\rho(x)$ for any $x\in\mathcal{V}\setminus\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$ and $\rho^{\prime}(\vec{x})=\vec{w}$) such that $\mathcal{M}\vDash_{\mathcal{M},\rho^{\prime}}\varphi$. An existentially constrained term is a triple $\langle X,s,\exists\vec{x}.\ \varphi\rangle$, written as $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, of a set $X$ of variables, a term $s$, and an existentially constraint $\exists\vec{x}.\ \varphi$, such that $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X\subseteq\mathcal{V}\mathsf{ar}(s)$, and $\mathcal{B}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\cap\mathcal{V}\mathsf{ar}(s)=\varnothing$. Variables in $X$ are called logical variables (of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$). An existentially constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is said to be satisfiable if $\exists\vec{x}.\ \varphi$ is satisfiable. Continued from the example above, $\exists x.\ x+y\geqslant 2$ is an example of existential constraint. We have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists x.\ x+y\geqslant 2)=\{y\}$ and $\mathcal{B}\mathcal{V}\mathsf{ar}(\exists x.\ x+y\geqslant 2)=\{x\}$. $\exists x,y.\ x+y\geqslant 2$ and $x+y\geqslant 2$ are also existential constraints; note that for the latter the empty sequence of variables is taken as $\vec{x}$, and we usually omit $\exists\varepsilon$. On the other hand, we exclude $\exists z.\ x+y\geqslant 2$ from the set of existential constraints, as $z\notin\{x,y\}$. Take $\rho$ as in the example above, and we have $\vDash_{\mathcal{M},\rho}\exists w.\ x+y\geqslant w$. $\Pi\{x,y\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358pt\exists w.\ x+y\geqslant w\hskip 0.85358pt]$ and $\Pi\{x,y\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358ptx\geqslant y\hskip 0.85358pt]$ are examples of existentially constrained terms. The $\Pi$-notation is intended to allow imposing value-only instantiations for variables not appearing in the constraint-part. For example, $\Pi\{x\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$ intends instantiations $\mathsf{sum}(n,t)$ for $n\in\mathbb{Z}$ and $t\in\mathcal{T}(\mathcal{F},\mathcal{V})$.
An existentially constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is said to be subsumed by an existentially constrained term $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, denoted by $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, if for all $X$-valued substitutions $\sigma$ with $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$ there exists a $Y$-valued substitution $\gamma$ with $\gamma\vDash_{\mathcal{M}}\exists\vec{y}.\ \psi$ such that $s\sigma=t\gamma$. Two existentially constrained terms $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ and $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ are said to be equivalent, denoted by $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\sim\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ if $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ and $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$. Continued from the example above, $\Pi\{x,y\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358ptx\geqslant y\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi\{x^{\prime},y^{\prime}\}.\ \mathsf{sum}(x^{\prime},y^{\prime})\penalty 10000\ [\hskip 0.85358pt\exists w.\ x^{\prime}+y^{\prime}\geqslant w\hskip 0.85358pt]$ and $\Pi\{x,y\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi\{x^{\prime}\}.\ \mathsf{sum}(x^{\prime},y^{\prime})\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$, both of which are not equivalent. We have $\Pi\{x,y\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358ptx=1\land y>x\hskip 0.85358pt]\sim\Pi\{y^{\prime}\}.\ \mathsf{sum}(1,y^{\prime})\penalty 10000\ [\hskip 0.85358pty^{\prime}\geqslant 2\hskip 0.85358pt]$ and $\Pi\{x,y\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358ptx\geqslant y\hskip 0.85358pt]\sim\Pi\{x^{\prime},y^{\prime}\}.\ \mathsf{sum}(x^{\prime},y^{\prime})\penalty 10000\ [\hskip 0.85358pt\exists w.\ w\geqslant 0\land x^{\prime}=y^{\prime}+w\hskip 0.85358pt]$. For various characterizations of equivalence, we refer to [17].
A constrained rewrite rule is a quadtuple consisting of a set $Z$ of theory variables, terms $\ell,r$ of a same sort and constraint $\pi$ such that $\mathcal{V}\mathsf{ar}(\varphi)\cup(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))\subseteq Z$, which is written as $\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$. A constrained rule $\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is said to be left-linear (left-value-free) if $\ell$ is linear (value-free, respectively). The original definition of constrained rewriting in [8, 4] is defined modulo equivalence $\sim$ of constrained terms, which makes analyses more complicated and its implementation mostly infeasible. On the other hand, most general constrained rewriting on existential constrained terms [18] has certain important advantages over the original definition of constrained rewriting, even though it is formulated only for left-linear constrained rewrite rules (up to now). As this paper is along the line of our previous paper [18], we again restrict ourselves to left-linear constrained rewrite rules.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term. Suppose that $\rho\colon\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is a left-linear constrained rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. We say $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$ if 1. $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, 2. $s|_{p}=\ell\gamma$, 3. $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and 4. $\vDash_{\mathcal{M}}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$, where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$. Suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$. Then we have a most general rewrite step $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ (or $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ with explicit $p$ and $\gamma$) where 1. $t=s[r\gamma]$, 2. $\psi=\varphi\land\pi\gamma$, 3. $\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\psi)\setminus\mathcal{V}\mathsf{ar}(t)$, and 4. $Y=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t))$. An LCTRS is a set of constrained rewrite rules that includes the set of calculation rules $\mathcal{R}_{\textsf{ca}}=\{\hskip 0.85358ptf(x_{1},\ldots,x_{n})\to y\penalty 10000\ [\hskip 0.85358pty=f(x_{1},\ldots,x_{n})\hskip 0.85358pt]\mid f\in\mathcal{F}_{\mathsf{th}}\setminus\mathcal{V}\mathsf{al}\hskip 0.85358pt\}$ (however, rules in $\mathcal{R}_{\textsf{ca}}$ will be not explicitly presented when we mention to $\mathcal{R}$ in examples). For an LCTRS $\mathcal{R}$, we have $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ if $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\rho\in\mathcal{R}$. The transitive and reflexive closure $\rightarrow_{\mathcal{R}}^{*}$ of $\rightarrow_{\mathcal{R}}$ is called the most general constrained reduction (of $\mathcal{R}$).
Example 2.1.
The constraint rewrite rules in Example˜1.1 in our formulation are presented as $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow 0\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$ and $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358ptx>0\hskip 0.85358pt]$, both of which are left-linear. Let $s$ be a constrained term $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]$. To perform a rewrite step from $s$, we take a renamed variant of the rule such as $\rho:\Pi\{\hskip 0.85358ptx^{\prime}\hskip 0.85358pt\}.\ \mathsf{sum}(x^{\prime})\rightarrow x^{\prime}+\mathsf{sum}(x^{\prime}+-1)\penalty 10000\ [\hskip 0.85358ptx^{\prime}>0\hskip 0.85358pt]$. Then $s$ has a $\rho$-redex at position $\epsilon$ using the substitution $\gamma=\{x^{\prime}\mapsto x\}$, as $\vDash_{\mathcal{M}}x>2\Rightarrow x>0$. From this we obtain a most general rewrite step $s\to^{\epsilon}_{\rho,\gamma}\Pi\{x\}.\ x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358ptx>2\land x>0\hskip 0.85358pt]$. In this case, the additional constraint $x>0$ is auxiliary, but for example, for the rewrite rule $\rho^{\prime}:\Pi\{\hskip 0.85358ptx^{\prime},y^{\prime}\hskip 0.85358pt\}.\ \mathsf{f}(x^{\prime})\rightarrow x^{\prime}+y^{\prime}\penalty 10000\ [\hskip 0.85358pty^{\prime}\geqslant x^{\prime}\hskip 0.85358pt]$, this addition is necessary, e.g. $\Pi\{x\}.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]\to_{\rho^{\prime}}\Pi\{y^{\prime}\}.\ x+y^{\prime}\penalty 10000\ [\hskip 0.85358ptx>2\land y^{\prime}>x\hskip 0.85358pt]$. Note that the extra variable $y^{\prime}$ in the rule remains not instantiated as the domain of the substitution equals to $\mathcal{V}\mathsf{ar}(\ell)$.
3 Partial Reduction of Constrained Terms
As explained in Example˜1.1, the existing notion of rewriting logically constrained terms is not able to capture situations in which a rewrite rule can be applied only to a part of the instantiations (here denoted as target constrained term). To reflect such a situation in this section, we introduce a novel notion called partial constrained reduction.
First, given a constrained rewrite rule $\rho$, we define the notion of a partial $\rho$-redex before we proceed to the definition of the partial rewrite steps.
Definition 3.1 (Partial $\rho$-Redex).
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term. Suppose that $\rho\colon\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is a left-linear constrained rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. We say $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a partial $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$ if the following is satisfied: 1. $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, 2. $s|_{p}=\ell\gamma$, 3. $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and 4. $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ is satisfiable in $\mathcal{M}$ where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$.
The difference to the definition of $\rho$-redex for a most general rewrite step lies in item 4. For a (non-partial) $\rho$-redex, we require the validity of the formula $(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$, whereas for a partial $\rho$-redex, we only require the satisfiability of the formula $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$. It follows that every (non-partial) $\rho$-redex is also a partial $\rho$-redex. To be precise, this fact is stated as follows.
Lemma 3.2.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term. Suppose that $\rho\colon\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is a left-linear constrained rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. Any $\rho$-redex of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ at position $p\in\mathcal{P}\mathsf{os}(s)$ using a substitution $\gamma$ is a partial $\rho$-redex at $p$ using $\gamma$.
Proof 3.3.
Suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$. Clearly, it suffices to show that item 4 in Definition˜3.1 follows. Since $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is satisfiable, $\exists\vec{x}.\ \varphi$ is satisfiable. Thus, there exists a valuation $\rho$ such that $\vDash_{\mathcal{M},\rho}\exists\vec{x}.\ \varphi$. By our assumption, $\mathcal{M}\vDash_{\rho}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$. Thus, $\vDash_{\mathcal{M},\rho}\exists\vec{z}.\ \pi\gamma$. Hence $\mathcal{M}\vDash_{\rho}(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ from which it follows that $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ is satisfiable.
If a satisfiable constrained term contains a partial $\rho$-redex then partial constrained reduction can be performed.
Definition 3.4 (Partial Constrained Reduction).
[…]
The only difference to the most general rewrite step is that the redex must be partial. Consequently, for the same $\rho$-redex, the reducts obtained by a most general rewrite step and by a partial rewrite step are identical.
It follows from the definitions of $\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}$ and $Y$ that $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ is again an existentially constrained term; this can be shown similarly to the case of the most general rewrite step [18].
Example 3.5.
[…]
$s\leadsto^{p}_{\rho,\gamma}\Pi\varnothing.\ 1+0\penalty 10000\ [\hskip 0.85358pt\exists w,y.\ 1\leqslant w\land w\leqslant 5\land y=w-1\land 0\geqslant y\hskip 0.85358pt].$
Note that $s$ is a normal form with respect to the most general rewrite step; this is due to the fact that $\vDash_{\mathcal{M}}(\exists w.\,1\leqslant w\land w\leqslant 5\land y=w-1)\Rightarrow(0\geqslant y)$ does not hold. An obvious witness is the valuation $\rho$ with $\rho(y)=1$.
We now compare partial rewrite steps with the most general rewrite steps of existentially constrained terms. First, we show that every most general rewrite step is also a partial rewrite step.
Theorem 3.6.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ be existentially constrained terms. If $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ then $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
Proof 3.7.
Suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. Then $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a $\rho$-redex at $p$ using $\gamma$, which is a partial $\rho$-redex by Lemma˜3.2. Hence, we have a partial rewrite step $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
As we demonstrated in Example˜3.5, partial rewrite steps may not be most general rewrite steps. Thus, the converse does not hold in general. However, we present a characterization of the converse.
Lemma 3.8.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ be existentially constrained terms. Suppose that $\rho\colon\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is a left-linear constrained rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. If $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ then $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma\hskip 0.85358pt]\to_{\rho}\cdot\sim\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$.
Proof 3.9.
Suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. Then, $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a partial $\rho$-redex at $p$ using $\gamma$. Thus, we have (1) $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, (2) $s|_{p}=\ell\gamma$, (3) $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and (4) $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ is satisfiable in $\mathcal{M}$.
By $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$ and condition (1) we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)\subseteq\{\gamma(x)\mid x\in\mathcal{V}\mathsf{ar}(\ell)\cap\mathcal{V}\mathsf{ar}(\pi)\}$. Then, since $\mathcal{V}\mathsf{ar}(\pi)\subseteq Z$ by the definition of rewrite rule, it follows that (4.5) $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)\subseteq X$ by condition (3).
By $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$, we have $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\cap\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\varnothing$ and $\mathcal{V}\mathsf{ar}(\varphi)\cap\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\varnothing$. Furthermore, by $\mathcal{V}\mathsf{ar}(\pi\gamma)\subseteq\mathcal{V}\mathsf{ar}(s)\cup\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}$, we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\pi\gamma)\cap\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}=\varnothing$. Hence, $\exists\vec{x}\mkern 2.0mu\vphantom{x},\vec{z}.\ \varphi\land\pi\gamma$ is a prenex normal form of $(\exists\vec{x}\mkern 2.0mu\vphantom{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$. It follows that (5) $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma)=\mathcal{F}\mathcal{V}\mathsf{ar}((\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma))$ and (6) $\vDash_{\mathcal{M}}(\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma)\Leftrightarrow((\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma))$.
By the definition of constrained terms, we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$. Hence, it follows from (4.5) and (5) that (7) $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma)\subseteq X$.
By condition (1), we have $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}\subseteq\mathcal{V}\mathsf{ar}(\pi\gamma)$. Then, since $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\subseteq\mathcal{V}\mathsf{ar}(\varphi)$, we have $\{\hskip 0.85358pt\vec{x},\vec{z}\hskip 0.85358pt\}\subseteq\mathcal{V}\mathsf{ar}(\varphi\land\pi\gamma)$. Also by $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\cap\mathcal{V}\mathsf{ar}(s)=\varnothing$ and $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$, we have $\{\hskip 0.85358pt\vec{x},\vec{z}\hskip 0.85358pt\}\cap\mathcal{V}\mathsf{ar}(s)=\varnothing$. Together with (7) we have that, $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma\hskip 0.85358pt]$ is an existentially constrained term.
From (6), we have $\vDash_{\mathcal{M}}(\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma)\Rightarrow(\exists\vec{x}.\ \varphi)$ and $\vDash_{\mathcal{M}}(\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$. The former gives $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$.
Consider a renamed variant $\rho^{\prime}:\Pi Z^{\prime}.\ \ell^{\prime}\rightarrow r^{\prime}\penalty 10000\ [\hskip 0.85358pt\pi^{\prime}\hskip 0.85358pt]$ of $\rho$ satisfying $\mathcal{V}\mathsf{ar}(\rho^{\prime})\cap\mathcal{V}\mathsf{ar}(s,\varphi\land\pi\gamma)=\varnothing$. Let $\delta$ be a renaming such that $\rho^{\prime}=\rho\delta$, and let $\gamma^{\prime}=\gamma\circ\delta^{-1}$. Then, we have $\ell^{\prime}\gamma^{\prime}=\ell\gamma$, $r^{\prime}\gamma^{\prime}=r\gamma$ and $\pi^{\prime}\gamma^{\prime}=\pi\gamma$. It also follows that $\vDash_{\mathcal{M}}(\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma)\Rightarrow(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi^{\prime}\gamma^{\prime})$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi^{\prime})\setminus\mathcal{V}\mathsf{ar}(\ell^{\prime})$. We obtain that $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma\hskip 0.85358pt]$ has a $\rho^{\prime}$-redex at $p$ using $\gamma^{\prime}$, and $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x},\vec{z}.\ \varphi\land\pi\gamma\hskip 0.85358pt]\to_{\rho}\Pi Y^{\prime}.\ t^{\prime}\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}\mkern 2.0mu\vphantom{y}^{\prime}.\ \psi^{\prime}\hskip 0.85358pt]$, where $t^{\prime}=s[r^{\prime}\gamma^{\prime}]_{p}$, $\psi^{\prime}=\varphi\land\pi\gamma\land\pi^{\prime}\gamma^{\prime}$, $\{\hskip 0.85358pt\vec{y}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\psi^{\prime})\setminus\mathcal{V}\mathsf{ar}(t^{\prime})$, and $Y^{\prime}=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t^{\prime}))$. Observe now that $t^{\prime}=s[r^{\prime}\gamma^{\prime}]_{p}=s[r\gamma]_{p}=t$, that $\vDash_{\mathcal{M}}\psi\Leftrightarrow\psi^{\prime}$ by $\psi=\varphi\land\pi\gamma$ and $\psi^{\prime}=\varphi\land\pi\gamma\land\pi^{\prime}\gamma^{\prime}=\varphi\land\pi\gamma\land\pi\gamma$, and that $\mathcal{V}\mathsf{ar}(\psi^{\prime})=\mathcal{V}\mathsf{ar}(\psi)$. Hence, $\Pi Y^{\prime}.\ t^{\prime}\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}\mkern 2.0mu\vphantom{y}^{\prime}.\ \psi^{\prime}\hskip 0.85358pt]\sim\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
The following is implied by LABEL:lem:from_weak_rewrite_step_to_mostgeneral_step.
Theorem 3.10.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ be existentially constrained terms. If $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ then $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\supset\\[-0.86108pt] \textstyle\sim\end{subarray}}\cdot\to_{\rho}\cdot\sim\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
4 Interpretations of Partial and Most General Constrained Reductions
It is natural to interpret an (existentially) constrained term as the set of instances of the term part that satisfy the constrained part. In this section, we provide interpretations of the most general rewrite steps and partial rewrite steps with respect to this interpretation of existentially constrained terms. This offers a more precise characterization of each type of rewrite step, allowing the two notions of constrained rewriting to be more clearly distinguished.
The idea that an existentially constrained term represents a set of terms (cf. [4, Section 2.3]) is formalized in our setting as follows.
Definition 4.1 (Interpretation of Existentially Constrained Terms).
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term. A term $t$ is said to be an instance of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ if there exists an $X$-valued substitution $\sigma$ such that $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$ and $t=s\sigma$. We denote the set of all instances of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ by $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, i.e.,
[…]
Note that by unfolding the definition of $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$, we have $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle=\{s\sigma\mid\sigma(X\cup\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi))\subseteq\mathcal{V}\mathsf{al},\vDash_{\mathcal{M}}(\exists\vec{x}.\ \varphi)\sigma\}$.
Example 4.2.
We have $\langle\!\langle\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,z)\penalty 10000\ [\hskip 0.85358pt\exists y.\ x=y\times 2\hskip 0.85358pt]\rangle\!\rangle=\{\mathsf{f}(n,w)\mid n\in\mathbb{Z}/2,w\in\mathcal{T}(\mathcal{F},\mathcal{V})\}$.
The following are some immediate properties of the interpretation.
Lemma 4.3.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term. Then,
  1. 1.
    $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is satisfiable iff $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle\neq\varnothing$, and
  2. 2.
    the set $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ is closed under substitutions.
Subsumption and equivalence of constrained terms are characterized in terms of their interpretation.
Lemma 4.4.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ be existentially constrained terms. Then,
  1. 1.
    $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ iff $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle\subseteq\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$, and
  2. 2.
    $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\sim\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ iff $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle=\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$.
Proof 4.5.
Clearly 2 immediately follows from 1 by the definition of equivalence. To show 1, let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ be existentially constrained terms. We start by proving the only if-direction. Assume that $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. To show $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle\subseteq\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$, take an element $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. Then, by definition, there exists an $X$-valued substitution $\sigma$ with $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$ and $u=s\sigma$. Then, by definition of $\mathrel${$\begin{subarray}{c}\textstyle\subset\\[-0.86108pt]$ $\sim\end{subarray}$}, there exists a $Y$-valued substitution $\gamma$ with $\gamma\vDash_{\mathcal{M}}\exists\vec{y}.\ \psi$ and $s\sigma=t\gamma$. It follows that $u=s\sigma=t\gamma\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$. For the if-direction we assume that $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle\subseteq\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$ holds. Fix an arbitrary $X$-valued substitution $\sigma$ with $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$. Then, as we have $s\sigma\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, by our assumption $s\sigma\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle=\{\hskip 0.85358ptt\gamma\mid\gamma(Y)\subseteq\mathcal{V}\mathsf{al},\gamma\vDash_{\mathcal{M}}\exists\vec{y}.\ \psi\hskip 0.85358pt\}$. Thus, $s\sigma=t\gamma$ for some $Y$-valued substitution $\gamma$ such that $\gamma\vDash_{\mathcal{M}}\exists\vec{y}.\ \psi$. Hence, we conclude $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
The following properties, which directly follow from the definition, are immediate corollaries of Lemma˜4.4:
  • $\mathrel${$\begin{subarray}{c}\textstyle\subset\\[-0.86108pt]$ $\sim\end{subarray}$} forms a quasi-order,
  • $\sim$ is an equivalence relation,
  • all unsatisfiable existentially constrained terms are equivalent, and
  • no satisfiable existentially constrained term is equivalent to any unsatisfiable existentially constrained term.
The interpretation of rewrite rules is defined analogously, by considering synchronized instances of both sides of each rule.
Definition 4.6 (Interpretation of Constrained Rules).
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear constrained rewrite rule. We define the interpretation $\langle\!\langle\rho\rangle\!\rangle$ of $\rho$ as: $\langle\!\langle\rho\rangle\!\rangle=\{\langle\ell\sigma,r\sigma\rangle\mid\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r),\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\}$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$. For an LCTRS $\mathcal{R}$, we define $\langle\!\langle\mathcal{R}\rangle\!\rangle=\bigcup_{\rho\in\mathcal{R}}\langle\!\langle\rho\rangle\!\rangle$.
Example 4.7.
Let $\rho:\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,y)\rightarrow y\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$ be a constrained rewrite rule. Then $\langle\!\langle\rho\rangle\!\rangle=\{\mathsf{f}(n,y)\to y\mid 0\geqslant n\in\mathbb{Z}\}$.
Note that, in contrast to the interpretation of existentially constrained terms, non-logical variables of rewrite rules are not instantiated in the interpretation of rewrite rules. We note also that, unlike the interpretation of existentially constrained terms, the non-logical variables in rewrite rules are not instantiated in the interpretation of rewrite rules.
We continue with characterizations of most general rewrite steps and partial rewrite steps in terms of their interpretation. We begin with a characterization of partial rewrite steps.
Lemma 4.8.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. Then,
  1. 1.
    there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some $v$,
  2. 2.
    if $w\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$ then there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}w$, and
  3. 3.
    for any $u,v$ such that $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ and $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$, we have $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$.
Proof 4.9.
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear constrained rewrite rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$ together with $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. We have (1) $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, (2) $s|_{p}=\ell\gamma$, (3) $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and (4) $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ is satisfiable where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$. Furthermore, we have (5) $t=s[r\gamma]_{p}$, (6) $\psi=\varphi\land\pi\gamma$, (7) $\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\psi)\setminus\mathcal{V}\mathsf{ar}(t)$, and (8) $Y=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t))$.
  1. 1.
[…]
    Let $u=s\hat{\delta}$ and let us show that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some $v$. For this, it suffices to show $\ell\gamma\hat{\delta}\to^{\epsilon}_{\langle\!\langle\rho\rangle\!\rangle}r\gamma\hat{\delta}$ as $u|_{p}=(s\hat{\delta})|_{p}=s|_{p}\hat{\delta}=\ell\gamma\hat{\delta}$. Let $\sigma=(\hat{\delta}\circ\gamma)|_{Z\cap\mathcal{V}\mathsf{ar}(\ell,r)}$. It remains to show that $\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$ and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
    For the former, by our choice of $\sigma$, it remains to show $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. Assume $z\in Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$. We distinguish two cases. If $z\in Z\cap\mathcal{V}\mathsf{ar}(\ell)$, then we have $\gamma(z)\in X$ by (3); thus, $\sigma(z)=\hat{\delta}(\gamma(z))\in\mathcal{V}\mathsf{al}$ by our choice of $\hat{\delta}$. Otherwise, we have $z\in Z\cap\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$. Then, $\sigma(z)=\hat{\delta}(\gamma(z))=\hat{\delta}(z)\in\mathcal{V}\mathsf{al}$ by $\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{D}\mathsf{om}(\gamma)=\varnothing$ and our choice of $\hat{\delta}$.
    To show $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$, let $z\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)$, i.e., $z\in\mathcal{V}\mathsf{ar}(\pi)\cap\mathcal{V}\mathsf{ar}(\ell,r)$. If $z\in\mathcal{V}\mathsf{ar}(\ell)$, then $z\in\mathcal{V}\mathsf{ar}(\ell)\cap X$ by $z\in\mathcal{V}\mathsf{ar}(\pi)$, and hence $\gamma(z)\in\mathcal{V}\mathsf{al}$. Thus, $\sigma(z)=\hat{\delta}(\gamma(z))\in\mathcal{V}\mathsf{al}$ follows. Otherwise $z\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$. Then $\gamma(z)=z$, and hence $\sigma(z)=\hat{\delta}(z)\in\mathcal{V}\mathsf{al}$ by our choice of $\hat{\delta}$. Thus, $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$ and it remains to show $\vDash_{\sigma}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$.
    By our choice of $\hat{\delta}$, we have $\vDash_{\hat{\delta}}\pi\gamma$. Thus, $\vDash_{[\![\gamma]\!]_{\hat{\delta}}}\pi$. Hence $\vDash_{[\![\gamma]\!]_{\hat{\delta}}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$. We now show $\gamma]\!]_{\hat{\delta}}(x)=\sigma(x)$ for all $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)$. From $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)$ it follows $x\in\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$, and hence $x\notin\mathcal{V}\mathsf{ar}(\ell)$. Thus, $\gamma(x)=x$. Hence $\gamma(x)]\!]_{\hat{\delta}}=[\![x]\!]_{\hat{\delta}}=\hat{\delta}(x)=\hat{\delta}(\gamma(x))=\sigma(x)$ follows.
    Hence we conclude the claim.
  2. 2.
    Suppose $w\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$. Then, by definition, (9) $w=t\delta$, (10) $\delta(Y)\subseteq\mathcal{V}\mathsf{al}$, and (11) $\delta\vDash_{\mathcal{M}}\exists\vec{y}.\ \psi$ for some substitution $\delta$. without loss of generality assume $\mathcal{D}\mathsf{om}(\delta)\subseteq\mathcal{V}\mathsf{ar}(t)$.
    As $\mathcal{V}\mathsf{ar}(t)\cap\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\varnothing$, one can extend $\delta$ to $\delta^{\prime}=\delta\cup\{\vec{y}\mapsto\vec{v}\}$ so that $\vDash_{\delta^{\prime}}\psi$ for some $\vec{v}\in\mathcal{V}\mathsf{al}^{*}$. We further extend $\delta^{\prime}$ to $\delta^{\prime\prime}=\delta^{\prime}\cup\{\vec{x}\mkern 2.0mu\vphantom{x}\mapsto\vec{w}\}$ where $\{\hskip 0.85358pt\vec{x}\mkern 2.0mu\vphantom{x}\hskip 0.85358pt\}=X\setminus\mathcal{V}\mathsf{ar}(t)$ by taking an arbitrary $\vec{w}\in\mathcal{V}\mathsf{al}^{*}$.
    We now show $s\delta^{\prime\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. Firstly, by $Y=(X\cap\mathcal{V}\mathsf{ar}(t))\cup(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))$ and (10), we have $X\cap\mathcal{V}\mathsf{ar}(t)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta)$. Thus, $\delta^{\prime\prime}(X)\subseteq\mathcal{V}\mathsf{al}$, by our choice of $\delta^{\prime\prime}$. It remains to show $\delta^{\prime\prime}\vDash\exists\vec{x}.\ \varphi$. We have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{y}.\ \psi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta)$ by (11). Hence, $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq\mathcal{V}\mathsf{ar}(\varphi)\subseteq\mathcal{V}\mathsf{ar}(\psi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta^{\prime})\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta^{\prime\prime})$. Also, by $\vDash_{\delta^{\prime}}\psi\penalty 10000\ (=\varphi\land\pi\gamma)$, we have $\vDash_{\delta^{\prime}}\exists\vec{x}.\ \varphi$. Thus, since $\delta^{\prime}(x)=\delta^{\prime\prime}(x)$ for $x\notin\{\vec{x}\}$, we have $\vDash_{\delta^{\prime\prime}}\exists\vec{x}.\ \varphi$. Hence $\delta^{\prime\prime}\vDash\exists\vec{x}.\ \varphi$ as claimed. Thus, $s\delta^{\prime\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$.
    Let $u=s\delta^{\prime\prime}$. Since $\delta(x)=\delta^{\prime\prime}(x)$ for all $x\in\mathcal{V}\mathsf{ar}(t)$, we have $u|_{p}=s|_{p}\delta^{\prime\prime}=\ell\gamma\delta^{\prime\prime}$ and $w=t\delta=t\delta^{\prime\prime}=s[r\gamma]_{p}\delta^{\prime\prime}=s\delta^{\prime\prime}[r\gamma\delta^{\prime\prime}]_{p}=u[r\gamma\delta^{\prime\prime}]_{p}$. Thus, to obtain $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}w$, it remains to show $\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$ and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$, where $\sigma$ is a substitution with $\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$ given by $\sigma(x)=\delta^{\prime\prime}(\gamma(x))$.
    For the former, it suffices to show $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$ by our choice of $\sigma$. Let $z\in Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$. If $z\in Z\cap\mathcal{V}\mathsf{ar}(\ell)$ then $\gamma(z)\in\mathcal{V}\mathsf{al}\cup X$. Then, by $\delta^{\prime\prime}(X)\subseteq\mathcal{V}\mathsf{al}$, it follows $\sigma(z)=\delta^{\prime\prime}(\gamma(z))\in\mathcal{V}\mathsf{al}$. If $z\in Z\cap\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$, then $z\in Y$ by the definition of rewrite steps, and hence $\sigma(z)=\delta^{\prime\prime}(\gamma(z))=\delta(z)\in\mathcal{V}\mathsf{al}$ by (10). Hence, all in all, $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$ holds.
    As $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$, we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. Furthermore, from $\vDash_{\delta^{\prime}}\psi\penalty 10000\ (=\varphi\land\pi\gamma)$ and $\mathcal{V}\mathsf{ar}(\pi\gamma)\cap\{\vec{x}\}=\varnothing$, it follows that $\vDash_{\delta^{\prime\prime}}\pi\gamma$. Hence, we have $\vDash_{[\![\gamma]\!]_{\delta^{\prime\prime}}}\pi$. By $\mathcal{D}\mathsf{om}(\gamma)\cap\mathcal{V}\mathsf{ar}(\pi)\subseteq\mathcal{V}\mathsf{ar}(\ell)\cap X$, we have $\gamma]\!]_{\delta^{\prime\prime}}(x)=\delta^{\prime\prime}(\gamma(x))=\sigma(x)$ for all $x\in\mathcal{V}\mathsf{ar}(\pi)$. Hence, $\vDash_{\sigma}\pi$. Thus, $\vDash_{\sigma}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$, and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ follows by $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. This concludes the claim.
  3. 3.
    Suppose $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ and $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$. Then $u=s\delta$ for substitution $\delta$ such that $\delta(X)\subseteq\mathcal{V}\mathsf{al}$ and $\delta\vDash\exists\vec{x}.\ \varphi$. without loss of generality suppose $\mathcal{D}\mathsf{om}(\delta)\subseteq\mathcal{V}\mathsf{ar}(s)$.
    From $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$, there exists a substitution $\sigma$ such that $u|_{p}=\ell\sigma$, $v=u[r\sigma]_{p}$, $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$ and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$. Since $u|_{p}=s|_{p}\delta=\ell\gamma\delta$, we have $\sigma(x)=\delta(\gamma(x))$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)$. We are now going to show $v=u[r\sigma]_{p}\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$.
    By $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$, we know $\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. Thus, for $x\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$, we have $\sigma(x)\in\mathcal{V}\mathsf{al}$. Thus, if $x=r|_{q_{x}}$ then $\sigma(x)=(r\sigma)|_{q_{x}}=v|_{p.q_{x}}$ holds. Thus, one can extend $\delta$ to $\delta^{\prime}$ by $\delta^{\prime}(x)=v|_{p.q_{x}}$ for $x\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$. Then, for $x\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$, we have $\sigma(x)=v|_{p.q_{x}}=\delta^{\prime}(x)=\delta^{\prime}(\gamma(x))$ and also clearly, $\delta(x)=\delta^{\prime}(x)$ holds also for $x\in\mathcal{V}\mathsf{ar}(\ell)$. Thus, we have $\sigma(x)=\delta^{\prime}(\gamma(x))$ for all $x\in\mathcal{V}\mathsf{ar}(\ell,r)$.
    Note that $\mathcal{V}\mathsf{ar}(s)\cap\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)=\varnothing$, we have $s\delta=s\delta^{\prime}$. Thus, $u[r\sigma]_{p}=s\delta[r\sigma]_{p}=s\delta^{\prime}[r\gamma\delta^{\prime}]_{p}=(s[r\gamma]_{p})\delta^{\prime}=t\delta^{\prime}$.
    Thus, it remains to to show $\delta^{\prime}(Y)\subseteq\mathcal{V}\mathsf{al}$ and $\delta^{\prime}\vDash\exists\vec{y}.\ \psi$, i.e., $\delta^{\prime}\vDash\exists\vec{y}.\ \varphi\land\pi\gamma$.
    Let $y\in Y=(X\cap\mathcal{V}\mathsf{ar}(t))\cup\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$. If $y\in X\cap\mathcal{V}\mathsf{ar}(t)\subseteq X$ then $\delta^{\prime}(y)=\delta(y)\in\mathcal{V}\mathsf{al}$ by $\delta(X)\subseteq\mathcal{V}\mathsf{al}$. If $y\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$, then $\delta^{\prime}(y)\in\mathcal{V}\mathsf{al}$ by our construction of $\delta^{\prime}$. Thus, $\delta^{\prime}(Y)\subseteq\mathcal{V}\mathsf{al}$ holds.
    As $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{y}.\ \psi)\subseteq Y$, $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{y}.\ \psi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta^{\prime})$ follows from $\delta^{\prime}(Y)\subseteq\mathcal{V}\mathsf{al}$. Thus, it remains to show $\vDash_{\delta^{\prime}}\exists\vec{y}.\ \psi$. From $\vDash_{\delta}\exists\vec{x}.\ \varphi$ we have $\vDash_{\delta^{\prime}}\exists\vec{x}.\ \varphi$.
    From $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$, we have $\vDash_{\delta^{\prime}\circ\gamma}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$. Thus, $\vDash_{\delta^{\prime}}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\gamma$. As $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$ and $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$, and $\mathcal{D}\mathsf{om}(\gamma)\cap\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\varnothing$. Also, by $\gamma(\mathcal{V}\mathsf{ar}(\ell))\subseteq\mathcal{V}\mathsf{ar}(s)$, and $\mathcal{R}\mathsf{an}(\gamma)\cap\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\varnothing$. Thus, $\vDash_{\delta^{\prime}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\gamma$. Hence $\vDash_{\delta^{\prime}}(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\gamma)$. From this, it follows $\vDash_{\delta^{\prime}}\exists\vec{x},\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \varphi\land\pi\gamma$. Since $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\cap\mathcal{V}\mathsf{ar}(t)=\varnothing$ and $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}\cap\mathcal{V}\mathsf{ar}(t)=\varnothing$, it follows $\vDash_{\delta^{\prime}}\exists\vec{y}.\ \varphi\land\pi\gamma$.
Lemma˜4.8 immediately yields the following characterization of partial constrained reductions in terms of interpretation.
Theorem 4.10 (Characterization by Interpretation for Partial Constrained Reduction).
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\stackrel{{\scriptstyle*}}{{\leadsto}}_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
  1. 1.
    For each $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$ there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ such that $u\stackrel{{\scriptstyle*}}{{\to}}_{\langle\!\langle\mathcal{R}\rangle\!\rangle}v$.
  2. 2.
    There exist $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ and $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$ such that $u\stackrel{{\scriptstyle*}}{{\to}}_{\langle\!\langle\mathcal{R}\rangle\!\rangle}v$.
Proof 4.11.
1 follows from item 2 of Lemma˜4.8. 2 follows from the items 1 and 3 of Lemma˜4.8.
Note that by Theorem˜3.6, the properties that hold for partial rewrite steps should hold also for most general rewrite steps. Thus, the properties presented in Theorem˜4.10 also hold for most general rewrite steps.
In fact, a property stronger than (2) holds for most general rewrite steps. To show this, the following lemma is used.
Lemma 4.12.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. For any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ there exists $v$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$.
Proof 4.13.
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear constrained rewrite rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$, and suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. Thus, we have (1) $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, (2) $s|_{p}=\ell\gamma$, (3) $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and (4) $\vDash_{\mathcal{M}}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$, where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$. Furthermore, we have (5) $t=s[r\gamma]_{p}$, (6) $\psi=\varphi\land\pi\gamma$, (7) $\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\psi)\setminus\mathcal{V}\mathsf{ar}(t)$, and (8) $Y=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t))$.
Let $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. Then, by definition, (9) $u=s\delta$, (10) $\delta(X)\subseteq\mathcal{V}\mathsf{al}$, and (11) $\delta\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$ for some substitution $\delta$ such that $\mathcal{D}\mathsf{om}(\delta)\subseteq\mathcal{V}\mathsf{ar}(s)$.
Now we define a substitution $\sigma$ as follows:
  • For $x\in\mathcal{V}\mathsf{ar}(\ell)$, let $\sigma(x)=\delta(\gamma(x))$.
  • For $x\in\mathcal{V}\mathsf{ar}(r)\setminus(\ell)$, we define $\sigma(x)$ as follows. For this, let $\{\hskip 0.85358pt\vec{w}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell)$. Let $\hat{\delta}$ be a valuation such that $\hat{\delta}(x)=\delta(x)$ for all $x\in X$. Then, by (11) and $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$, we have $\vDash_{\hat{\delta}}\exists\vec{x}.\ \varphi$. Hence, by (4), we know $\vDash_{\hat{\delta}}(\exists\vec{z}.\ \pi\gamma)$. From this, it follows $\vDash_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\gamma$ for some $\vec{v}\in\mathcal{V}\mathsf{al}^{*}$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$ and $\hat{\delta}[\vec{w}\mapsto\vec{v}]$ denotes the update of the valuation $\hat{\delta}$ by $\vec{w}\mapsto\vec{v}$. Then, we take $\sigma(\vec{w})=\vec{v}$.
  • For $x\in\mathcal{V}\setminus\mathcal{V}\mathsf{ar}(\ell,r)$, we take an arbitrary $\sigma(x)$.
Here, for the later use, let $\hat{\delta}^{\prime}=\hat{\delta}[\vec{w}\mapsto\vec{v}]$. Clearly, since $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\cap\{\hskip 0.85358pt\vec{w}\hskip 0.85358pt\}=\varnothing$, we have $\vDash_{\hat{\delta}^{\prime}}\exists\vec{x}.\ \varphi$. and hence (12) $\vDash_{\hat{\delta}^{\prime}}(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\gamma)$ holds.
From (2) and (9), it follows that $u|_{p}=s\delta|_{p}=s|_{p}\delta=\ell\gamma\delta=\ell\sigma$ by our choice of $\sigma$ for $x\in\mathcal{V}\mathsf{ar}(\ell)$. To obtain $v$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$, it suffices to show $\ell\sigma\to_{\langle\!\langle\rho\rangle\!\rangle}r\sigma$. For this, we take $\sigma^{\prime}=\sigma|_{Z\cap\mathcal{V}\mathsf{ar}(\ell,r)}$ and show $\mathcal{D}\mathsf{om}(\sigma^{\prime})=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma^{\prime})=Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$ and $\sigma^{\prime}\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$.
The former follows from $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma^{\prime})$ by our choice of $\sigma^{\prime}$. Let $z\in Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$. If $z\in Z\cap\mathcal{V}\mathsf{ar}(\ell)$ then $\gamma(z)\in\mathcal{V}\mathsf{al}\cup X$ by (3), and thus, $\sigma^{\prime}(z)=\delta(\gamma(z))\in\mathcal{V}\mathsf{al}$ by (10). Otherwise $z\in(Z\cap\mathcal{V}\mathsf{ar}(r))\setminus(Z\cap\mathcal{V}\mathsf{ar}(\ell))\subseteq\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell)$, and hence $\sigma^{\prime}(z)\in\mathcal{V}\mathsf{al}$ by our choice of $\sigma^{\prime}$ for $z\in\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell)$. Thus, $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma^{\prime})$.
We proceed to show the latter. We first show $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma^{\prime})$. This follows from the former because $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)=\mathcal{V}\mathsf{ar}(\pi)\setminus(\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r))=\mathcal{V}\mathsf{ar}(\pi)\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$.
It remains to show $\vDash_{\hat{\sigma}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ for the valuation $\hat{\sigma}$ such that $\hat{\sigma}(z)=\sigma^{\prime}(z)$ for all $z\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)$. From $\mathcal{R}\mathsf{an}(\gamma)\subseteq\mathcal{V}\mathsf{ar}(s)$, $\mathcal{R}\mathsf{an}(\gamma)\cap\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\varnothing$. We also have $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$ and $\mathcal{V}\mathsf{ar}(\ell)\cap\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\varnothing$, and hence $\mathcal{D}\mathsf{om}(\gamma)\cap\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\varnothing$. Thus, it follows from $\vDash_{\hat{\delta}^{\prime}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\gamma$ that $\vDash_{\hat{\delta}^{\prime}}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\gamma$. Hence, we have $\vDash_{[\![\gamma]\!]_{\hat{\delta}^{\prime}}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$.
We now show $\gamma]\!]_{\hat{\delta}^{\prime}}(z)=\hat{\sigma}(z)$ for all $z\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)=\mathcal{V}\mathsf{ar}(\pi)\cap\mathcal{V}\mathsf{ar}(\ell,r)$. From this, our claim $\vDash_{\hat{\sigma}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ clearly follows. We distinguish two cases.
  • Suppose $z\in\mathcal{V}\mathsf{ar}(\pi)\cap\mathcal{V}\mathsf{ar}(\ell)$. Then, $\gamma]\!]_{\hat{\delta}^{\prime}}(z)=[\![\gamma]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}(z)=[\![\gamma(z)]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}=[\![\gamma(z)]\!]_{\hat{\delta}}=\hat{\delta}(\gamma(z))=\delta(\gamma(z))=\hat{\sigma}(z)$, as $\gamma(z)\in X\cup\mathcal{V}\mathsf{al}$.
  • Otherwise, let $z\in\mathcal{V}\mathsf{ar}(\pi)\cap(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))$. Then, $\gamma]\!]_{\hat{\delta}^{\prime}}(z)=[\![\gamma]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}(z)=[\![\gamma(z)]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}=[\![z]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}=\{\vec{w}\mapsto\vec{v}\}(z)=\hat{\sigma}(z)$.
This concludes the claim.
We now arrive at a characterization of most general constrained reductions via interpretations.
Theorem 4.14 (Characterization by Interpretation for Most General Constrained Reduction).
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\stackrel{{\scriptstyle*}}{{\to}}_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. Then,
  1. 1.
    for each $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ there exists $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$ such that $u\stackrel{{\scriptstyle*}}{{\to}}_{\langle\!\langle\mathcal{R}\rangle\!\rangle}v$, and
  2. 2.
    for each $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$ there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ such that $u\stackrel{{\scriptstyle*}}{{\to}}_{\langle\!\langle\mathcal{R}\rangle\!\rangle}v$.
Proof 4.15.
1 follows from Lemma˜4.12, and 2 follows from Theorems˜3.6 and 4.10.
Note that the property of Theorem˜4.14 1 together with consistency implies the property of Theorem˜4.10 2. On the other hand, the property of Theorem˜4.14 1 does not hold for partial rewrite steps, as shown in the following example.
Example 4.16.
Consider Example˜3.5. We have
[…]
for $\rho:\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow 0\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$. Let $s=1+\mathsf{sum}(1)$. Then $s\in\langle\!\langle\Pi\{\hskip 0.85358pty\hskip 0.85358pt\}.\ 1+\mathsf{sum}(y)\penalty 10000\ [\hskip 0.85358pt\exists w.\ 1\leqslant w\land w\leqslant 5\land y=w-1\hskip 0.85358pt]\rangle\!\rangle$. However, we have $\langle\!\langle\Pi\varnothing.\ 1+0\penalty 10000\ [\hskip 0.85358pt\exists w,y.\ 1\leqslant w\land w\leqslant 5\land y=w-1\land 0\geqslant y\hskip 0.85358pt]\rangle\!\rangle=\{1+0\}$, and $1+\mathsf{sum}(1)\to_{\rho}1+0$ does not hold. In fact, $s$ is a normal form with respect to the rule $\rho$.
Thus, the property of Theorem˜4.14 1 discriminates most general rewrite steps from partial rewrite steps. Then, a natural question arise: Is there any property that holds for partial rewrite steps, but not for most general rewrite steps?
Example 4.17.
Consider Example˜4.16. We have $\langle\!\langle\Pi\{\hskip 0.85358pty\hskip 0.85358pt\}.\ 1+\mathsf{sum}(y)\penalty 10000\ [\hskip 0.85358pt\exists w.\ 1\leqslant w\land w\leqslant 5\land y=w-1\hskip 0.85358pt]\rangle\!\rangle=\{1+\mathsf{sum}(n)\mid 0\leqslant n\leqslant 4\}$. The reason for the partial rewrite step by the rule $\rho:\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow 0\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$ is the existence of $1+\mathsf{sum}(0)$ in this set. On the other hand, $s^{\prime}=\Pi\{\hskip 0.85358pty\hskip 0.85358pt\}.\ 1+\mathsf{sum}(w)\penalty 10000\ [\hskip 0.85358pt1\leqslant w\land w\leqslant 5\hskip 0.85358pt]$ is a normal form with respect to the rule $\rho$. In this case, we have $\langle\!\langle s^{\prime}\rangle\!\rangle=\{1+\mathsf{sum}(n)\mid 1\leqslant n\leqslant 5\}$, whose members are all normal forms with respect to the rule $\rho$. This motivates us to characterize normal forms of partial write steps by the absence of reducible terms in interpretations. However, the interpretation so far is not useful for this. Consider the rule $\rho^{\prime}:\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,y)\rightarrow y\penalty 10000\ [\hskip 0.85358pt0=x\hskip 0.85358pt]$ and the existentially constrained term $t=\Pi\{\hskip 0.85358ptw\hskip 0.85358pt\}.\ \mathsf{f}(w,z)\penalty 10000\ [\hskip 0.85358pt1\leqslant w\land w\leqslant 5\hskip 0.85358pt]$. Then, although $t$ is normal with respect to the rule $\rho^{\prime}$, $\langle\!\langle t\rangle\!\rangle$ includes the term $\mathsf{f}(1,\mathsf{f}(0,z))$, which is reducible by the rule $\rho^{\prime}$. For a different illustration, consider the following rewrite rule $\rho^{\prime\prime}:\Pi\varnothing.\mathsf{f}(\mathsf{a})\to\mathsf{b}[\mathsf{true}]$ and the constrained term $\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$. On the one hand $\mathsf{f}(\mathsf{a})\in\langle\!\langle\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle$ is reducible, while on the other hand $\mathsf{f}(\mathsf{b})\in\langle\!\langle\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle$ is a normal form. Thus the standard interpretation can hardly be used to distinguish normal form instances from reducible instances.
In the following sections, we give a characterization of partial rewrite steps which does not hold for most general rewrite steps, by introducing another kind of interpretations.
5 Value Interpretation of Existentially Constrained Terms
The standard interpretation of existentially constrained terms in LABEL:sec:Interpretations_of_Weak_and_Most_GeneralReductions is inadequate for characterizing partial rewrite steps by their normal forms. In this section, we introduce an alternative interpretation of constrained terms. The idea is that logical variables are instantiated by values, as before, but for non-logical variables we only allow renaming instead of proper instantiation.
Definition 5.1 (Value Interpretation of Constrained Terms).
A term $t$ is said to be a value instance of an existentially constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ if $t=s\sigma$ for some substitution $\sigma$ such that 1. $\sigma(X)\subseteq\mathcal{V}\mathsf{al}$, 2. $\sigma(\mathcal{V}\mathsf{ar}(s)\setminus X)\subseteq\mathcal{V}$, 3. $x\neq y$ implies $\sigma(x)\neq\sigma(y)$ for any variable $x,y\in\mathcal{V}\mathsf{ar}(s)\setminus X$, and 4. $\vDash_{\mathcal{M},\sigma}\exists\vec{x}.\ \varphi$. We denote the set of all value instances of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ by $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, which is called the value interpretation of the constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$.
Note here that, as $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$, it follows from the items 1 and 3 that $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$. The reason to allow renaming of non-logical variables for value interpretations is to enable it to characterize equivalence ($\sim$) between (existentially) constrained terms (see Theorem˜5.8).
Example 5.2.
We have $\langle\!\langle\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,z)\penalty 10000\ [\hskip 0.85358pt\exists y.\ x=y\times 2\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}=\{\mathsf{f}(n,w)\mid n\in\mathbb{Z}/2,w\in\mathcal{V}\}$.
Some immediate properties of value instantiations follow.
Lemma 5.3.
1. An existentially constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is satisfiable iff $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}\neq\varnothing$. 2. The set $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ is closed under variable renaming. 3. $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}\subseteq\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. 4. $\bigcup_{\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}}\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle=\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. TA: What do you mean by $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$???
We focus now on less obvious properties of value instantiations. We first clarify notable relations between value instantiations and instantiations.
Lemma 5.4.
For any existentially constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, we have $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle=\{t\theta\mid t\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}},\mbox{$\theta$ is a substitution}\}$.
Proof 5.5.
($\subseteq$) Suppose $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. Then there exists a substitution $\sigma$ such that $u=s\sigma$, $\sigma(X)\subseteq\mathcal{V}\mathsf{al}$ and $\sigma\vDash\exists\vec{x}.\ \varphi$. Take $\sigma^{\prime}:=\sigma|_{X}$ and $\theta:=\sigma|_{\overline{X}}$. Then, (1) $\sigma^{\prime}(X)=\sigma(X)\subseteq\mathcal{V}\mathsf{al}$, (2) since $\sigma^{\prime}(x)=x\in\mathcal{V}$ for $x\in\mathcal{V}\mathsf{ar}(s)\setminus X$, $\sigma^{\prime}(\mathcal{V}\mathsf{ar}(s)\setminus X)\subseteq\mathcal{V}$ holds, and (3) since $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$ and $\sigma^{\prime}(x)=\sigma(x)$ for any $x\in X$, $\sigma^{\prime}\vDash\exists\vec{x}.\ \varphi$ holds. Thus, it follows that $s\sigma^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. Finally, since $\sigma=\theta\circ\sigma^{\prime}$, we obtain $u=s\sigma=s\sigma^{\prime}\theta\in\{t\theta\mid t\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, $\theta$ is a substitution}\}. ($\supseteq$) Suppose $u=t\theta$ for some substitution $\theta$ and $t\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. Then, $t=s\sigma$ for some $\sigma$ such that (1) $\sigma(X)\subseteq\mathcal{V}\mathsf{al}$, (2) $\sigma(\mathcal{V}\mathsf{ar}(s)\setminus X)\subseteq\mathcal{V}$, (3) $\sigma(x)\neq\sigma(y)$ for any $x,y\in\mathcal{V}\mathsf{ar}(s)\setminus X$ with $x\neq y$, and (4) $\sigma\vDash\exists\vec{x}.\ \varphi$. Take $\sigma^{\prime}:=\theta\circ\sigma$. Then, $\sigma^{\prime}(x)=\theta(\sigma(x))=\sigma(x)\in\mathcal{V}\mathsf{al}$ for any $x\in X$. Thus, $\sigma^{\prime}(X)=\sigma(X)\subseteq\mathcal{V}\mathsf{al}$. Also, $\sigma^{\prime}\vDash\exists\vec{x}.\ \varphi$ follows by $\sigma\vDash\exists\vec{x}.\ \varphi$ and $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$. Thus, $s\sigma^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. Finally, the claim follows as $u=t\theta=s\sigma\theta=s\sigma^{\prime}$.
From now on we write $s\leqslant t$ if $s\theta=t$.
Lemma 5.6.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term. Then, $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ is the set of terms minimal with respect to $\leqslant$ in $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$.
Proof 5.7.
The claim follows from (1) $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}\subseteq\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, and (2) if $u,v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ with $u<v$, then $v\notin\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. (1) is immediate from Lemma˜5.3 3. To show (2), suppose $u,v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ with $u<v$, i.e., there exist substitutions $\sigma,\theta$ such that (3) $u=s\sigma$, (4) $\sigma(X)\subseteq\mathcal{V}\mathsf{al}$, (5) $\sigma\vDash\exists\vec{x}.\ \varphi$, (6) $v=u\theta$, and (7) $v\not\leqslant u$. Note that from (3)–(5) we obtain $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, and thus $v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ follows by Lemma˜4.3 2 and 6. Now we suppose $v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ and derive a contradiction. Let $s=s[x_{1},\ldots,x_{m},y_{1},\ldots,y_{n}]_{p_{1},\ldots,p_{m},q_{1},\ldots,q_{n}}$ with $\mathcal{P}\mathsf{os}_{X}(s)=\{p_{1},\ldots,p_{m}\}$ and $\mathcal{P}\mathsf{os}_{\mathcal{V}\setminus X}(s)=\{q_{1},\ldots,q_{n}\}$. Then, by (3) and (4), $u=s[v_{1},\ldots,v_{m},y_{1}\sigma,\ldots,y_{n}\sigma]$ with $v_{1},\ldots,v_{m}\in\mathcal{V}\mathsf{al}$. By (6), $v=u\theta=s[v_{1},\ldots,v_{m},y_{1}\sigma\theta,\ldots,y_{n}\sigma\theta]$. From our assumption that $v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, it follows that the sequence of terms $\langle y_{1}\sigma\theta,\ldots,y_{n}\sigma\theta\rangle$ is a sequence of variables which is a renaming of $\langle y_{1},\ldots,y_{n}\rangle$. Hence, $v\leqslant u$. This is a contradiction to (7).
Lemma˜5.6 enables us to show that equivalence of constrained terms is characterized by value instantiation.
Theorem 5.8.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ be existentially constrained terms. Then, $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\sim\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ iff $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}=\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
Proof 5.9.
We show that $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle=\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$ iff $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}=\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. The if-part follows from Lemma˜5.4, while the only-if part follows from Lemma˜5.6. The full claim then follows from Lemma˜4.4 2.
Contrast to the equivalence $\sim$ of existentially constraint terms, the value interpretation does not reflect subsumption relation on it, i.e., the following is not correct: $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ iff $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle\subseteq\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle$. This is witnessed in the following example.
Example 5.10.
Consider the two existential constrained terms $\Pi\{x\}.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$ and $\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$. Then on the one hand we have $\Pi\{x\}.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$. On the other hand, $\langle\!\langle\Pi\{x\}.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}\subseteq\langle\!\langle\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ does not hold. It gets more obvious by $\langle\!\langle\Pi\{x\}.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}=\{\mathsf{f}(v)\mid v\in\mathcal{V}\mathsf{al}\}\neq\{\mathsf{f}(y)\mid y\in\mathcal{V}\}=\langle\!\langle\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
Note that if one defines the value interpretation without allowing to rename non-logical variables, then Theorem˜5.8 does not hold. Consider $\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\sim\Pi\varnothing.\ \mathsf{f}(y)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$ for which we obtain $\langle\!\langle\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}=\{\mathsf{f}(x)\}\neq\{\mathsf{f}(y)\}=\langle\!\langle\Pi\varnothing.\ \mathsf{f}(y)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
6 Characterization of Partial and Most General Constrained Reductions by Value Instantiation
In this section, we provide a characterization of partial constrained reductions and most general constrained reductions with respect to their normal forms, expressed in terms of value interpretation. In contrast to the characterizations in Section˜4, this characterization distinguishes partial rewrite steps from most general rewrite steps. Firstly, we extend Lemma˜4.8 to value interpretation.
Lemma 6.1.
Suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. Then all of the following holds:
  1. 1.
    There exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some $v$.
  2. 2.
    If $w\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ then there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}w$.
  3. 3.
    If $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ and $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$, then $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
Proof 6.2.
  1. 1.
    The specific substitution $\hat{\delta}$ in the proof of Lemma˜4.8 satisfies $\hat{\delta}(x)=x$ for $x\in\mathcal{V}\mathsf{ar}(s)\setminus X$, as $(\mathcal{V}\mathsf{ar}(\pi\gamma)\cup\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\setminus X)\cap\mathcal{V}\mathsf{ar}(s)=\varnothing$. Thus, it follows that $u=s\hat{\delta}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
  2. 2.
    In the proof of Lemma˜4.8, assume moreover that (10’) $\delta(\mathcal{V}\mathsf{ar}(t)\setminus Y)\subseteq\mathcal{V}$ and (10”) for any $\in\mathcal{V}\mathsf{ar}(t)\setminus Y, xyx\neq y$ implies $\delta(x)\neq\delta(y)$. From the definitions of $\delta^{\prime}$ and $\delta^{\prime\prime}$, clearly they also satisfy these properties. Since $(\mathcal{V}\mathsf{ar}(t)\setminus Y)\subseteq(\mathcal{V}\mathsf{ar}(s)\setminus X)$, it follows that $u=s\delta^{\prime\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
  3. 3.
    In the proof of Lemma˜4.8, assume moreover that $\delta(\mathcal{V}\mathsf{ar}(s)\setminus X)\subseteq\mathcal{V}$ and for any $\in\mathcal{V}\mathsf{ar}(s)\setminus X, xyx\neq y$ implies $\delta(x)\neq\delta(y)$. To show $t\delta^{\prime}\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, it remains to show (1) $\delta^{\prime}(\mathcal{V}\mathsf{ar}(t)\setminus Y)\subseteq\mathcal{V}$ and (2) for any $\in\mathcal{V}\mathsf{ar}(t)\setminus Y, xyx\neq y$ implies $\delta(x)\neq\delta(y)$. They follow as $\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\subseteq Y$, $\mathcal{V}\mathsf{ar}(t)\subseteq\mathcal{V}\mathsf{ar}(s)\cup\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$, and $(\mathcal{V}\mathsf{ar}(t)\setminus Y)\subseteq(\mathcal{V}\mathsf{ar}(s)\setminus X)$.
Now we proceed with a characterization in the reverse direction, that is, when a partial rewrite step from a constrained term is guaranteed by the rewrite step of its instantiation.
Lemma 6.3.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term. Suppose there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some position $p$ and a term $v$. Then, there exists an existentially constrained term $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ such that $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ and $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
Proof 6.4.
Let $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. Then, $u=s\mu$ for a substitution $\mu$ such that (1) $\mu(X)\subseteq\mathcal{V}\mathsf{al}$, (2) $\mu(\mathcal{V}\mathsf{ar}(s)\setminus X)\subseteq\mathcal{V}$, and (3) $\mu\vDash\exists\vec{x}.\ \varphi$. Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear constrained rewrite rule without loss of generality such that $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$, and suppose that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$. More precisely, suppose that $\sigma$ is a substitution such that (4) $\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$ and (5) $\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$, (6) $u|_{p}=\ell\sigma\delta$, and (7) $v=u[r\sigma\delta]_{p}$ so that $\ell\sigma\to r\sigma\in\langle\!\langle\rho\rangle\!\rangle$ and $u\to^{p}_{\ell\sigma\to r\sigma}v$.
We first show $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\gamma$. From (1) and (2), we have $\mu(\mathcal{V}\mathsf{ar}(s))\subseteq\mathcal{V}\mathsf{al}\cup\mathcal{V}$. Thus, $\mathcal{P}\mathsf{os}(s|_{p})=\mathcal{P}\mathsf{os}(s\mu|_{p})=\mathcal{P}\mathsf{os}(\ell\sigma)$. Hence, as $\ell$ is linear and value-free, there exists a substitution $\gamma$ such that $\ell\gamma=s|_{p}$ and $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$. Using (6), we now have $\ell\gamma\mu=s|_{p}\mu=s\mu|_{p}=u|_{p}=\ell\sigma\delta$, i.e., (8) $\mu(\gamma(x))=\delta(\sigma(x))$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)$.
Let $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$. Then (9) $\mu(\gamma(x))=\delta(\sigma(x))=\sigma(x)\in\mathcal{V}\mathsf{al}$ by (4) and (8). Thus $\gamma(x)\in\mathcal{V}\mathsf{al}\cup\mathcal{V}$. Moreover, $\gamma(x)\in\mathcal{V}\mathsf{al}\cup\mathcal{V}\mathsf{ar}(s)$ by $\ell\gamma=s|_{p}$. Hence $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$; for, if $\gamma(x)\in\mathcal{V}\mathsf{ar}(s)\setminus X$ then $\mu(\gamma(x))\in\mathcal{V}$ by (2), which contradicts $\mu(\gamma(x))\in\mathcal{V}\mathsf{al}$. Thus, (10) $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$.
Let $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$. From (5) and $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}\subseteq\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}$, we have $\vDash_{\sigma}\exists\vec{z}.\ \pi$. As $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi)=\mathcal{V}\mathsf{ar}(\ell)\cap Z$, we have $\vDash_{\mu\circ\gamma}\exists\vec{z}.\ \pi$ by (9). Thus, $\vDash_{\mu}(\exists\vec{z}.\ \pi)\gamma$. By $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}\cap\mathcal{D}\mathsf{om}(\gamma)=\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}\cap\mathcal{R}\mathsf{an}(\gamma)=\varnothing$, we have $\vDash_{\mu}\exists\vec{z}.\ \pi\gamma$. Thus, together with (3), we have (11) $\vDash_{\mu}(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$.
All in all, we know that $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, $s|_{p}=\ell\gamma$, $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$ by (10), and $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ is satisfiable by (11). Thus, we have $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, where (15) $t=s[r\gamma]_{p}$, (16) $\psi=\varphi\land\pi\gamma$, (17) $\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\psi)\setminus\mathcal{V}\mathsf{ar}(t)$, and (18) $Y=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t))$. It remains to show $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
Let $\mu^{\prime}$ by a substitution given by $\mu^{\prime}(x)=\mu(x)$ for $x\in\mathcal{V}\mathsf{ar}(s)\cup X$, $\mu^{\prime}(x)=\sigma(x)$ for each $x\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$, and $\mu^{\prime}(x)=x$ otherwise. Clearly, (19) $s\mu=s\mu^{\prime}$.
Let $y\in\mathcal{V}\mathsf{ar}(r)$. If $y\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$ then $\mu^{\prime}(\gamma(y))=\mu^{\prime}(y)=\sigma(y)=\delta(\sigma(y))$, as $\sigma(y)\in\mathcal{V}\mathsf{al}$ by (4). Otherwise $y\in\mathcal{V}\mathsf{ar}(\ell)$. Then, $\mu(\gamma(y))=\delta(\sigma(y))$ by (8). Also, as $\gamma(y)$ is a subterm of $s$, $\mu^{\prime}(\gamma(y))=\mu(\gamma(y))$ by (19). Hence $\mu^{\prime}(\gamma(y))=\delta(\sigma(y))$. Thus, $\mu^{\prime}(\gamma(y))=\delta(\sigma(y))$ for all $y\in\mathcal{V}\mathsf{ar}(r)$, and so (20) $r\gamma\mu^{\prime}=r\sigma\delta$.
Now, using (19) and (20), we have (21) $v=u[r\sigma\delta]_{p}=s\mu[r\sigma\delta]_{p}=s\mu^{\prime}[r\gamma\mu^{\prime}]_{p}=s[r\gamma]_{p}\mu^{\prime}=t\mu^{\prime}$.
Let $y\in Y=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t))$. If $y\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)$, then $y\in\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(r)\cap Z$, and hence, $\mu^{\prime}(y)=\sigma(y)\in\mathcal{V}\mathsf{al}$ by (4). If $y\in X\cap\mathcal{V}\mathsf{ar}(t)$ then $\mu^{\prime}(y)=\mu(y)$ by $y\in X$, and hence it follow from $y\in X$ that $\mu^{\prime}(y)=\mu(y)\in\mathcal{V}\mathsf{al}$ by (1). Thus, (22) $\mu^{\prime}(Y)\subseteq\mathcal{V}\mathsf{al}$ holds.
Let $y\in\mathcal{V}\mathsf{ar}(t)\setminus Y$. Then $y\in\mathcal{V}\mathsf{ar}(t)\setminus(\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t)))$. Thus, $y\in\mathcal{V}\mathsf{ar}(t)\setminus(\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup X)$. Clearly, if $y\notin\mathcal{V}\mathsf{ar}(s)\cup X$, then $\mu^{\prime}(x)=x\in\mathcal{V}$. If $y\in\mathcal{V}\mathsf{ar}(s)\cup X$, then $y\in\mathcal{V}\mathsf{ar}(s)\setminus X$, and hence $\mu^{\prime}(y)=\mu(y)\in\mathcal{V}$ by (2). Thus, (23) $\mu^{\prime}(\mathcal{V}\mathsf{ar}(t)\setminus Y)\subseteq\mathcal{V}$ holds.
We now show $\vDash_{\mu^{\prime}}\exists\vec{y}.\ \psi$. From (11), we have $\vDash_{\mu}\exists\vec{x}.\ \varphi$ and $\vDash_{\mu}\exists\vec{z}.\ \pi\gamma$. As $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$, we have $\vDash_{\mu^{\prime}}\exists\vec{x}.\ \varphi$. Also, by $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)\subseteq\mathcal{V}\mathsf{ar}(s)$ we have $\vDash_{\mu^{\prime}}\exists\vec{z}.\ \pi\gamma$. Since $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\cap\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\varnothing$, one can extend $\mu^{\prime}$ to $\mu^{\prime\prime}$ so that $\vDash_{\mu^{\prime\prime}}\varphi\land\pi\gamma$. Since $\{\hskip 0.85358pt\vec{x},\vec{z}\hskip 0.85358pt\}\subseteq\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}$, it follows (24) $\vDash_{\mu^{\prime}}\exists\vec{y}.\ \psi$.
Thus, from (21)–(24), we conclude $v\in\langle\!\langle\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.
Theorem 6.5 (Normal Forms of Partial Constrained Reduction).
Suppose $\mathcal{R}$ is a left-linear and left-value-free LCTRS. Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term. Then, $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is a normal form with respect to $\leadsto$$_{\mathcal{R}}$ iff any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ is a normal form with respect to $\to_{\langle\!\langle\mathcal{R}\rangle\!\rangle}$.
Proof 6.6.
[…]
($\Longrightarrow$) We proceed with a proof by contradiction. Suppose there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ such that $u\to_{\langle\!\langle\mathcal{R}\rangle\!\rangle}^{p}v$ for some $v$. Then, by the definition of value instantiation, $u=s\mu$ for some $X$-valued substitution $\mu$ such that $\mathcal{D}\mathsf{om}(\mu)=X$ and $\mu\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$. By $u\to_{\langle\!\langle\mathcal{R}\rangle\!\rangle}^{p}v$, we have $\langle u|_{p},v|_{p}\rangle\in\langle\!\langle\rho\rangle\!\rangle$ for some $\rho\in\mathcal{R}$. However, it follows from Lemma˜6.3 that $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
Note that the theorem above can be reformulated as follows. The following statements are equivalent: (i) $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, and (ii) there exists some $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ such that $u\to_{\langle\!\langle\mathcal{R}\rangle\!\rangle}v$ for some $v$.
Example 6.7.
The ($\Longrightarrow$)-direction of Theorem˜6.5 does not hold if we replace the value interpretation $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ with the standard interpretation $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. To see this, let us revisit Example˜4.17. Consider the rule $\rho^{\prime}:\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,y)\rightarrow y\penalty 10000\ [\hskip 0.85358pt0=x\hskip 0.85358pt]$ and the existentially constrained term $t=\Pi\{\hskip 0.85358ptw\hskip 0.85358pt\}.\ \mathsf{f}(w,z)\penalty 10000\ [\hskip 0.85358pt1\leqslant w\land w\leqslant 5\hskip 0.85358pt]$. Then $\mathsf{f}(1,\mathsf{f}(0,z))\in\langle\!\langle t\rangle\!\rangle$ and $\mathsf{f}(1,\mathsf{f}(0,z))\to^{2}_{\langle\!\langle\rho^{\prime}\rangle\!\rangle}\mathsf{f}(1,z)$ by the rule $\mathsf{f}(\mathsf{0},y)\to y\in\langle\!\langle\rho^{\prime}\rangle\!\rangle$. However, there does not exists an $s$ such that $t\leadsto^{2}_{\rho^{\prime}}s$. Consider the rule $\rho^{\prime\prime}:\Pi\varnothing.\mathsf{f}(\mathsf{a})\to\mathsf{b}[\mathsf{true}]$ and the constrained term $t^{\prime}=\Pi\varnothing.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$. Then $\mathsf{f}(\mathsf{a})\in\langle\!\langle t^{\prime}\rangle\!\rangle$ and $\mathsf{f}(\mathsf{a})\to^{\epsilon}_{\langle\!\langle\rho^{\prime\prime}\rangle\!\rangle}\mathsf{b}$ by using the rule $\mathsf{f}(\mathsf{a})\to\mathsf{b}\in\langle\!\langle\rho^{\prime\prime}\rangle\!\rangle$. However, there does not exist an $s^{\prime}$ such that $t^{\prime}\leadsto^{\epsilon}_{\rho^{\prime\prime}}s^{\prime}$.
Under the view of Theorems˜4.10 and 4.14, one might conjecture that if one replaces “any uu” in the statement of Theorem˜6.5 by “there exists uu”, then the normal form of constrained terms with respect to most general constrained reduction can be obtained. However, this is not the case. Consider $\mathcal{R}$ in Example˜1.1 and the constrained term $\mathsf{sum}(x)\penalty 10000\ [0\leqslant x\land x\leqslant 1]$. Then $\langle\!\langle\mathsf{sum}(x)\penalty 10000\ [0\leqslant x\land x\leqslant 1]\rangle\!\rangle_{\mathsf{v}}=\{\mathsf{sum}(0),\mathsf{sum}(1)\}$ and we have $\mathsf{sum}(0)\to_{\langle\!\langle\mathcal{R}\rangle\!\rangle}0$ and $\mathsf{sum}(1)\to_{\langle\!\langle\mathcal{R}\rangle\!\rangle}1+\mathsf{sum}(1+-1)$. However, $\mathsf{sum}(x)\penalty 10000\ [0\leqslant x\land x\leqslant 1]$ is a normal form with respect to $\to_{\mathcal{R}}$.
We are now going to show a variation of Theorem˜6.5 for the most general rewrite step, by modifying the form of the necessary and sufficient condition of the theorem.
Lemma 6.8.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term and $\rho$ a left-value-free and left-linear rewrite rule. Let $p\in\mathcal{P}\mathsf{os}(s)$. Suppose that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ there exists $v$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$. Then $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
Proof 6.9.
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear and left-value-free constrained rewrite rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. Suppose that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ there exists $v$ such that $\langle u|_{p},v|_{p}\rangle\in\{\langle\ell\sigma,r\sigma\rangle\mid Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma),\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\}$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
It suffices to show that $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a $\rho$-redex at $p$ by some $\gamma$. That is, we have to show there exists a substitution $\gamma$ such that (i) $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, (ii) $s|_{p}=\ell\gamma$, (iii) $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and (iv) $\vDash_{\mathcal{M}}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$, where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$.
First, we show there exists $s|_{p}$ matches with $\ell$, that is, $s|_{p}=\ell\gamma$ for some substitution $\gamma$.
By the satisfiability of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, there exists $\delta$ such that $\delta\vDash\exists\vec{x}.\ \varphi$. Then $u=s\delta\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. Note here that (0) $\mathcal{P}\mathsf{os}(u)=\mathcal{P}\mathsf{os}(s\delta)=\mathcal{P}\mathsf{os}(s)$ and $\delta(x)\in\mathcal{V}\mathsf{al}$ for all $x\in X$ and $\delta(y)\in\mathcal{V}$ for all $y\in\mathcal{V}\setminus X$ by the definition of value instantiation. By our assumption, there exists a substitution $\sigma$ such that (1) $u|_{p}=\ell\sigma$, (2) $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$, and (3) $\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
Let $q\in\mathcal{P}\mathsf{os}_{\mathcal{F}}(\ell)$. Then $\ell_{q}=\ell\sigma|_{q}=u|_{p.q}=s\delta_{p.q}$. If $p.q\in\mathcal{P}\mathsf{os}_{\mathcal{F}}(s)$, then $\ell(q)=s\delta(p.q)=s(p.q)$. Otherwise $p.q\in\mathcal{P}\mathsf{os}_{\mathcal{V}}(s)$, as $\mathcal{P}\mathsf{os}(s)=\mathcal{P}\mathsf{os}(s\delta)$. Since $q\in\mathcal{P}\mathsf{os}_{\mathcal{F}}(\ell)$, it follows from $\ell(q)=s\delta(p.q)$ that $s\delta(p.q)\in\mathcal{V}\mathsf{al}$. However, this contradicts our assumption that $\ell$ is value-free. Thus, we have shown $s(p.q)=\ell(q)$ for all $q\in\mathcal{P}\mathsf{os}_{\mathcal{F}}(\ell)$. Hence, there exists a substitution $\gamma$ such that $s|_{p}=\ell\gamma$ by the left-linearity of $\rho$. without loss of generality one can assume $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$. Thus, we have shown (i) and (ii).
Now we have $\ell\gamma\delta=(s|_{p})\delta=s\delta|_{p}=u|_{p}=\ell\sigma$. Let $z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$ and $q$ be a position such that $\ell(q)=z$. Then we have $\sigma(z)\in\mathcal{V}\mathsf{al}$ by (2). Thus, $\delta(\gamma(z))=\ell\gamma\delta_{q}=\ell\sigma|_{q}=\sigma(z)\in\mathcal{V}\mathsf{al}$. Then $\gamma(z)\in\mathcal{V}\mathsf{al}\cup\mathcal{V}$. If $\gamma(z)=x\in\mathcal{V}$, then $\delta(x)\in\mathcal{V}\mathsf{al}$, and hence $x\in X$ by (0). Thus, $\gamma(z)\in\mathcal{V}\mathsf{al}\cup X$. Therefore, it holds that $\gamma(z)\in\mathcal{V}\mathsf{al}\cup X$ for any $z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$. Thus, we have shown (iii).
It remains to show (iv). For this, suppose $\xi$ is a valuation such that $\vDash_{\mathcal{M},\xi}\exists\vec{x}.\ \varphi$. Take a substitution $\hat{\xi}=\xi|_{X}$. Then $\hat{\xi}$ is $X$-valued and $\hat{\xi}\vDash\exists\vec{x}.\ \varphi$ by $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$. Also, as $\{\vec{x}\}\cap X=\varnothing$, we have ${\exists\vec{x}.\ \varphi\hat{\xi}}]\!]_{\mathcal{M}}=\mathsf{true}$. Let $u^{\prime}=s\hat{\xi}$. Then, $u^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. Thus, by our assumption, there exists a substitution $\sigma^{\prime}$ such that (1’) $u^{\prime}|_{p}=\ell\sigma^{\prime}$, (2’) $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma^{\prime})$, and (3’) $\sigma^{\prime}\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
Let $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)$, i.e., $x\in\mathcal{V}\mathsf{ar}(\pi\gamma)\setminus(\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell))$. Then $x\in\{\gamma(z)\in\mathcal{V}\mid z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z\}$. Thus, let $q,z$ be such that $z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, $x=\gamma(z)$, and $\ell|_{q}=z$. Then $\sigma^{\prime}(z)\in\mathcal{V}\mathsf{al}$ by (2’), and hence using (1’), we have $u^{\prime}|_{q}=\ell\sigma^{\prime}|_{q}=(\ell|_{q})\sigma^{\prime}=\sigma^{\prime}(z)\in\mathcal{V}\mathsf{al}$. Then, as $s\hat{\xi}=u^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, we have $s|_{q}\in X\cup\mathcal{V}\mathsf{al}$. Then, $\hat{\xi}(x)=\hat{\xi}(\gamma(z))=\ell|_{q}\gamma\hat{\xi}=(\ell\gamma|_{q})\hat{\xi}=(s|_{q})\hat{\xi}\in\mathcal{V}\mathsf{al}$, as $\hat{\xi}=\xi|_{X}$. Thus, for any $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)$, we have $\hat{\xi}(x)\in\mathcal{V}\mathsf{al}$. Furthermore, $\hat{\xi}(\gamma(z))=\hat{\xi}(x)=(s|_{q})\hat{\xi}=u^{\prime}|_{q}=\sigma^{\prime}(z)$. Thus, $\vDash_{\mathcal{M},\hat{\xi}}\exists\vec{z}.\ \pi\gamma$ by (3’). Hence, $\vDash_{\mathcal{M},\xi}\exists\vec{z}.\ \pi\gamma$. Thus, it follows that $\vDash_{\mathcal{M},\xi}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$ for any valuation $\xi$, and therefore (iv) holds.
Now it follows that Lemma˜4.12 can be strengthened with the converse direction by additional condition that the rewrite rule is left-value-free.
Corollary 6.10.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term and $\rho$ a left-linear and left-value-free rewrite rule. Let $p\in\mathcal{P}\mathsf{os}(s)$. Then the following statements are equivalent:
  1. 1.
    for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ there exists $v$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$, and
  2. 2.
    $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
Proof 6.11.
(2 $\Longrightarrow$ 1) follows by Lemma˜4.12. (1 $\Longrightarrow$ 2) follows from Lemma˜6.8 and Lemma˜5.3 3.
Now we arrive at a characterization of normal forms with respect to most general constrained reduction in terms of value interpretation. This is similar to the notion of normal form of constrained terms in [15, Definition 1].
Theorem 6.12 (Normal Forms of Most General Constrained Reduction).
Suppose $\mathcal{R}$ is a left-linear and left-value-free LCTRS. Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term. Then, $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is a normal form with respect to $\to_{\mathcal{R}}$ iff for any $\rho\in\mathcal{R}$ and position $p\in\mathcal{P}\mathsf{os}(s)$, there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for no $v$.
Proof 6.13.
[…]
($\Longrightarrow$) We prove the contraposition. Suppose there exist $\rho\in\mathcal{R}$ and $p\in\mathcal{P}\mathsf{os}(s)$ such that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some $v$. Then by Lemma˜6.8, $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
Note that the theorem above can be reformulated as follows. The following statements are equivalent: (i) $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, and (ii) there exists $\rho\in\mathcal{R}$ and position $p\in\mathcal{P}\mathsf{os}(s)$ such that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ there exists $v$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$. In contrast to the characterization of partial constrained reduction (Theorem˜6.5), our characterization of most general constrained reduction refers to reducibility of all $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ with respect to a fixed position and a fixed rewrite rule.
Example 6.14.
Let us recall the LCTRS $\mathcal{R}$ in Example˜1.1:
[…]
Consider the rewrite steps starting from the constrained term $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]$. Note that we have $\langle\!\langle\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}=\{\mathsf{sum}(0),\ldots,\mathsf{sum}(4)\}$, $\langle\!\langle\rho_{1}\rangle\!\rangle=\{\mathsf{sum}(n)\to 0\mid 0\geqslant n\in\mathbb{Z}\}$, and $\langle\!\langle\rho_{2}\rangle\!\rangle=\{\mathsf{sum}(n)\to n+\mathsf{sum}(n+-1)\mid 0<n\in\mathbb{Z}\}$. We further have $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]\leadsto^{\epsilon}_{\rho_{1}}\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ 0\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\land 0\geqslant x\hskip 0.85358pt]$, and $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]\leadsto^{\epsilon}_{\rho_{2}}\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\land 0<x\hskip 0.85358pt]$. By Theorem˜6.5, the former is characterized by $\mathsf{sum}(0)\to^{\epsilon}_{\langle\!\langle\rho\rangle\!\rangle}0$ and the latter by $\mathsf{sum}(4)\to^{\epsilon}_{\langle\!\langle\rho\rangle\!\rangle}4+\mathsf{sum}(4+-1)$. Also none of the TRSs $\langle\!\langle\rho_{1}\rangle\!\rangle,\langle\!\langle\rho_{2}\rangle\!\rangle$ can reduce all of the terms $\mathsf{sum}(0),\ldots,\mathsf{sum}(4)$. By Theorem˜6.12, we obtain that $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]$ is a normal form with respect to $\to_{\mathcal{R}}$.
7 Related Work
Partial constrained reduction implicitly appears in various forms in the applications and the analyses of LCTRSs and of a related formalism of rewriting with built-in data structures.
Critical Pair Analysis
Critical pair analysis is a standard technique for checking confluence of LCTRSs. Many approaches rely on closure conditions for constrained critical pairs. Splitting a constrained critical pair [15] partitions its set of instances into subsets, each of which may admit a rewrite step. The goal is that, after splitting, a concrete rewrite step becomes applicable and satisfies the closing condition, which corresponds to a partial rewrite step on the original constrained critical pair. Partial rewrite steps can also be used to determine whether such a split is possible, as exploited to some extent in tools like crest [15].
Partial rewrite steps can also be used to detect non-confluence in LCTRSs, since it suffices that an instance of a critical pair diverges. In this case, explicit splitting is not required: a partial step could directly expose the subset of instances responsible for divergence. Tools like crest exploit this principle to efficiently find counterexamples without actual partitioning.
Constrained Rewriting Induction
An equation is said to be an inductive theorem of an equational theory when it is valid in the initial algebra of the theory; in other words, when all ground instantiations of the equation are equational consequences of the theory (e.g. [5]). Rewriting induction, that has been initiated in [12], is an approach to verify inductive theorems based on rewriting techniques. The notion of inductive theorems and the rewriting induction approach have been incorporated to the LCTRS formalism in [4, 10], where the latter has been called “constrained rewriting induction.” A key inference rule in rewriting induction is the rule Expand, which consists of a collection of narrowing steps applied at an appropriate position in an equation using all rules applicable at that position. A narrowing step is a rewrite step that may involve a minimal instantiation of the target term, replacing matching with unification. In constrained rewriting induction, instead of using only pure narrowing steps, a combination of narrowing steps and partial constrained rewrite steps is employed. This approach collects the reducts of all possible rewrite steps for minimal instantiations, while potentially adding constraints to the constrained equation at an appropriate position. In addition, although not covered in [4], partial constrained reduction may be useful to strengthen the Disprove rule for refuting a conjecture in constrained rewriting induction, similarly to the part on critical pair analysis above.
Constrained Narrowing
“Constrained Narrowing” [3, 6] narrows both, instances of non-logical variables and constraints for logical variables: To apply a constrained rewrite rule to a constrained term, non-logical variables of the constrained term are minimally instantiated, and the constraint of the rule is added to the constraint of the constrained term. Thus, constrained narrowing can be considered as “Narrowing” + “Partial Rewriting”.
Rewriting Modulo SMT
“Rewriting Modulo SMT” [13] is a sophisticated and intricate rewriting formalism (in a broad sense) that is based on a framework and terminology different from that of LCTRSs. Nevertheless, it shares key features with LCTRSs, such as the ability to handle built-in data structures and the use of constrained rewrite rules and rewriting on constrained terms. Although precise comparisons between the two formalisms are beyond the scope of this paper, roughly speaking, this formalism is designed to model state transition systems for more application-oriented specifications. In contrast to LCTRSs, it includes extended features such as order-sorted signatures and rewriting modulo equations, as well as restrictions like rewriting only at the root position, focusing on initial models, and a hierarchical signature for term sorts over theory sorts. We would like to highlight two noteworthy correspondences between results in [13] and ours. First, the basic rewrite relation (referred to as the symbolic rewrite relation in [13]) corresponds more closely to partial constrained rewriting rather than to most general constrained rewriting. Second, papers establish soundness and completeness results with respect to the standard interpretations. Readers may compare the soundness results (Lemma˜4.8 and Lemma˜6.1 versus Theorem 1 of [13]) and completeness results (Lemma˜6.3 and Example˜6.7 versus Theorem 2 of [13]). Overall, our results help to clarify the differences between the rewriting modulo SMT and LCTRS formalisms and highlight the respective advantages of each.
8 Conclusion
In this paper, we have introduced a novel notion of partial constrained rewriting for existentially constrained terms. The key difference between the definition of most general constrained rewrite steps and that of partial constrained rewrite steps lies in the definition of redex—validity is needed for most general rewriting while satisfiability suffices for partial rewriting. We have shown that most general rewrite steps are partial, while partial rewrite steps can be simulated by the composition of subsumption, most general rewrite steps and equivalence. We have further developed rigorous characterizations of partial and most general constrained reductions through value and standard interpretations. We have shown that while the standard interpretation captures some behavior of partial and most general rewrite steps, it is insufficient to distinguish partial rewrite steps from most general rewrite steps. By introducing the notion of value interpretation, we fully describe normal forms of both types of constrained reductions, including successfully highlighting subtle differences between partial and most general constrained reductions. We have also revisited immature ideas of partial rewriting that appeared in the literature; we anticipate that our results not only clarify the theoretical structure of logically constrained rewriting but also lay the groundwork for further exploration of (non-)reachability, (non-)joinability and (non-)convertibility verification on logically constrained terms, that appear in many applications and analyses of LCTRSs.
We focus in this paper on constrained rewriting with left-linear constrained rewrite rules. Despite some complications, we anticipate that one can extend partial constrained rewriting to non-left-linear rules. We leave this extension as future work. Another question is, whether partial constrained rewriting and equivalence commute. We expect this to be the case, but solving this problem also remains as future work.
Appendix A Omitted Proofs